### v1

# NeSyGeo: A Neuro-Symbolic Framework for Multimodal Geometric Reasoning Data Generation

Abstract
Obtaining large-scale, high-quality data with reasoning paths is crucial for improving the geometric reasoning capabilities of multi-modal large language models (MLLMs). However, existing data generation methods, whether based on predefined templates or constrained symbolic provers, inevitably face diversity and numerical generalization limitations. To address these limitations, we propose NeSyGeo, a novel neuro-symbolic framework for generating geometric reasoning data. First, we propose a domain-specific language grounded in the entity–relation–constraint paradigm to comprehensively represent all components of plane geometry, along with generative actions defined within this symbolic space. We then design a symbolic–visual–text pipeline that synthesizes symbolic sequences, maps them to corresponding visual and textual representations, and generates diverse question–answer (Q&A) pairs using large language models (LLMs). To the best of our knowledge, we are the first to propose a neuro-symbolic approach in generating multimodal reasoning data. Based on this framework, we construct NeSyGeo-CoT and NeSyGeo-Caption datasets, containing 100k samples, and release a new benchmark NeSyGeo-Test for evaluating geometric reasoning abilities in MLLMs. Experiments demonstrate that the proposal significantly and consistently improves the performance of multiple MLLMs under both reinforcement and supervised fine-tuning. With only 4k samples and two epochs of reinforcement fine-tuning, base models achieve improvements of up to +15.8% on MathVision, +8.4% on MathVerse, and +7.3% on GeoQA. Notably, a 4B model can be improved to outperform an 8B model from the same series on geometric reasoning tasks.
1 Introduction
Improving the visual reasoning capabilities of MLLMs has garnered significant attention recently  [17, 2, 1, 14, 11, 33, 25, 28, 15], with models like InternVL [5] and the QwenVL series [27, 3] demonstrating significant enhancements in visual-semantic comprehension through their multimodal capabilities. Among various visual reasoning tasks, geometric mathematical reasoning is crucial for evaluating the reasoning performance of MLLMs [31, 30], as it requires a deep integration of spatial perception, symbolic understanding, and logical deduction. To enhance such reasoning abilities, existing approaches [37, 38, 39] primarily rely on fine-tuning base models using reinforcement learning (RL) or supervised fine-tuning (SFT) on specialized geometric reasoning datasets. These methods depend heavily on the availability of large-scale, high-quality geometric reasoning data, which is often costly and time-consuming to construct manually. Therefore, automatic data generation for geometric reasoning has emerged as a promising and actively explored direction, aiming to alleviate data scarcity and further improve the reasoning abilities of MLLMs.
Existing approaches for generating datasets in geometric tasks can be broadly classified into four categories. Text augmentation methods like G-LLaVA [9] primarily mutate the conditions of
Figure 1: Performance comparison of different MLLMs and LLMs with and without image input in several geometry datasets. The minimal or negligible drops observed upon image removal in GeoQA and R-CoT raise concerns regarding the utilization of visual information for geometric reasoning.
existing datasets through equivalent condition transformation and numerical scaling. However, this approach fails to address the scalability of image generation. Template-based methods [7, 37, 12], use predefined geometric templates with fixed topologies, simplifying synthesis but constraining diversity by reducing the geometric space to limited combinations. Solver-based methods [10, 32] inspired by symbolic prover AlphaGeometry [23], leverage formal languages for synthesis but lack metric details (e.g., angles, lengths, areas), restricting multimodal data to descriptive annotations and limiting numerical reasoning applications. Tool-based methods attempt to generate codes for tools like GeoGebra or MATLAB via LLMs. However, even advanced models struggle to ensure correctness with ambiguous natural language instructions and complex geometric spaces. In summary, existing methods grapple with issues of image scalability, limited geometric diversity, a lack of precise numerical information, and challenges in ensuring the reliability of generated content.
Beyond the challenges in data synthesis methodologies, current geometric reasoning datasets present several limitations that impede the advancement of MLLMs. A primary limitation stems from the often inadequate quality and low resolution of the provided images. Such inputs frequently fall below the optimal requirements of visual encoders [20, 16, 17], hindering the extraction of crucial fine-grained visual features and discriminative information essential for robust multimodal reasoning. Furthermore, our analysis reveals notable information redundancy between the textual and visual modalities in many current datasets. As shown in Figure 1, our comparative experiments demonstrate minimal or negligible accuracy drops upon image removal. This finding emphasizes the urgent need for a dataset that effectively separates textual and visual information and provides high-quality images to promote MLLMs’ visual perception and logical reasoning performance.
To address these challenges, we propose NeSyGeo, a neuro-symbolic framework for synthesising high-quality multimodal geometric reasoning datasets. NeSyGeo integrates three components: 1) A formal geometric symbolic space defined by a domain-specific language (DSL), capturing primitive entities (points, lines, circles), topological relations (parallelism, incidence, perpendicularity), and metric constraints (angles, lengths), enabling diverse geometric configurations via systematic sampling within constrained parametric bounds. 2) A bidirectional conversion engine that transforms symbolic constructs into decoupled modalities, producing annotated vector graphics paired with concise textual axioms. 3) A causal Q&A pairs and theorem-grounded Chain-of-Thought (CoT) sequences generator that effectively merges neural reasoning with symbolic verification. To our knowledge, we are the first to develop a neuro-symbolic framework for producing multimodal reasoning data.
Our framework enhances the diversity and validity of generated geometric reasoning data while effectively mitigating information redundancy and the underutilization of visual signals during training. Specifically, the comprehensive Geo-DSL and its expansive symbolic synthesis action-space promote diverse and well-grounded image generation. Meanwhile, our CoT sequence generator, powered by LLMs’ strong reasoning and language capabilities, conducts a backwards search across the geometric space to construct Q&A pairs, thereby enriching textual diversity. The unique identification of geometric elements via our symbolic language and dedicated conversion engine ensures visual validity. In parallel, a bidirectional cross-validation process using expert LLMs ensures textual validity. By strategically distributing complementary information across image and text modalities, our approach
Figure 2: Comparison of dataset characteristics synthesized by our method and other popular synthesis approaches. “High Resolution” denotes average image pixels exceeding 336$\times$336. “Symbolic Form” refers to the symbolic meta-information associated with the image. “Classification of Elements” signifies categorization by geometric elements. “Visual Understanding” represents the mitigation of image-text redundancy for stronger visual grounding in reasoning. More specific examples of different methods are in Appendix A.
encourages MLLMs to actively engage with visual information when solving problems, enhancing their abilities to perceive visually and effectively utilize images.
Leveraging the NeSyGeo pipeline, we construct two training datasets, NeSyGeo-Caption and NeSyGeo-CoT, comprising 100k samples. NeSyGeo-Caption aims to improve the perceptual understanding of geometric elements, while NeSyGeo-CoT primarily focuses on enhancing logical reasoning. The key characteristics of our dataset compared to other popular multimodal geometric datasets are presented in Figure 2. Additionally, we develop an evaluation set, NeSyGeo-Test, with 2668 Q&A pairs, enabling a thorough assessment of the geometric reasoning capabilities of mainstream MLLMs. We conducted an extensive and comprehensive evaluation of the geometric reasoning capabilities of current mainstream open-source and closed-source models, with details presented in Appendix 8. Notably, our training dataset consistently and efficiently enhances the geometric reasoning performance of MLLMs across multiple benchmarks. With only 4k samples and two epochs of RL training, base models achieve performance improvements of up to +15.8% on MathVision, +8.4% on MathVerse, and +7.3% on GeoQA. Moreover, InternVL2.5-4B can be improved to outperform the 8B model in the same series on geometric reasoning tasks.
In summary, our contributions are as follows:
  • We propose NeSyGeo, a novel framework for geometric reasoning data generation, featuring a Geo-DSL for symbolic synthesis, a conversion engine for image and text generation, and an LLM-driven generator for Q&A pairs with CoT. NeSyGeo ensures validity through rigorous symbolic definitions and diversity via varied actions and neural searching.
  • Using our framework, we synthesize the NeSyGeo-Caption and NeSyGeo-CoT training datasets with 100k high-quality samples, alongside a comprehensive geometric task evaluation set NeSyGeo-Test. These datasets are characterized by their diversity, rigor, and balanced distribution of information across image and text modalities.
  • We demonstrate significant performance improvements on several MLLMs across multiple benchmarks using both RL and SFT training methods with our training sets, validating the effectiveness of our framework and the high quality of our datasets.
2 Related Works
2.1 Geometric Problem-Solving
Early approaches to geometric reasoning predominantly relied on symbolic solvers that used formal languages to tackle the tasks. For instance, Inter-GPS [18] and PGDP [35] employed symbolic methods by manually crafting reasoning rules and symbolic representations for geometric entities. These systems typically transform visual input into symbolic forms through instance segmentation and apply theorem search to derive solutions. However, these methods lacked scalability due to their dependence on manually designed rules. Their inability to generalize beyond specific problem types further limited their universality and effectiveness across diverse geometric challenges.
Figure 3: The overview of our neuro-symbolic data generation framework. The framework comprises three steps: In the first step, we get a symbolic language sequence in a limited symbolic action space. In the second step, the conversion engine parses the Geo-DSL sequence and translates it back to natural language and visual image without losing soundness. In the third step, we employ LLMs to take a reverse search and forward validation process to get final Q&A pairs with CoT.
The advent of MLLMs has shifted the paradigm toward data-driven geometric reasoning, leveraging their robust reasoning capabilities. Recent advancements include GeoDRL [19]and GeoGen [13]. Despite these developments, geometric reasoning poses significant challenges for MLLMs, requiring seamless integration of image perception, geometric knowledge, and multi-step reasoning. GeoSense [29] identifies the identification and application of geometric principles as a persistent bottleneck. Similarly, GeoEval [34] reveals that current MLLMs exhibit significantly low accuracy when facing more challenging geometric problems. MathVerse [36] further highlights MLLMs’ over-reliance on textual information, underscoring the critical need for balanced multimodal datasets to enhance cross-modal reasoning capabilities.
2.2 Multimodal Geometry Datasets
Large-scale, high-quality datasets are essential for enhancing the performance of MLLMs in solving geometric problems. Early datasets such as GeoS [21] (186 problems), Geometry3k [18] (3000 problems) and GeoQA [4] (4998 problems) utilized human manual annotation. Their datasets are thus limited to a small scale. With the development of MLLMs, datasets of greater magnitude have become essential. To address this, numerous efforts have shifted toward automatic data generation.
G-LLaVA [9] rephrased questions from GeoQA and Geometry3k to create 115,000 Q&A pairs, but failed to enhance image variety. Template-based methods [37, 7] typically rely on 10–20 predefined geometric figures, limiting the diversity of the generated images. AlphaGeometry [23], a notable work that combines symbolic solvers for geometric proofs, employs a symbolic language definition. Yet, due to the absence of numerical attributes such as angle measures and segment lengths in its geometric space, attempts to automatically generate datasets using the AlphaGeometry framework [10, 32] are confined to caption datasets, failing to produce the numerical Q&A pairs critical for current MLLMs training. In contrast to prior approaches, our method pioneers a neuro-symbolic framework, being the first to integrate the precision of symbolic definition with the diversity of neural search for generating multimodal reasoning data.
3 Methods
To address the urgent need for large-scale and high-quality multimodal datasets in MLLMs for geometric reasoning, we propose NeSyGeo, a novel three-stage data generation pipeline. The pipeline is built upon Geo-DSL, a symbolic DSL designed to represent most elements in plane geometry space concisely and wholly. Its entity-relation-constraint structure allows any element to
Table 1: Examples of our Geo-DSL and Corresponding Natural Language. Geo-DSL is defined by an entity-relation-constraint framework, encompassing 13 point types, 7 line types, 3 angle types, and 14 shape types in plane geometry. See Appendix H for complete Geo-DSL definitions.
Type Geo-DSL Language Natural Language
Shape $Triangle(A,B,C)=(x,y,\alpha)$ Triangle $ABC$ has $AB=x$, $BC=y$, $\angle B=\alpha$
$Circle(O)=(x)$ Circle $O$ has radius $x$
Point $Foot(D,A,Line(B,C))$ $D$ is the foot of the perpendicular from $A$ to $BC$
$Intersection(E,Line(A,B),Line(C,D))$ $E$ is the intersection of line $AB$ and line $CD$
Line $Para(Line(A,B),Line(C,D),x)$ Line $AB$ is parallel to $CD$, $AB=x$
Angle $Angle(P,Q,R)=\alpha$ $\angle PQR=\alpha$
be defined via a single statement, while its expressive power ensures comprehensive coverage of all geometric elements and values.
NeSyGeo’s generation process unfolds in three distinct stages: First, a symbolic generator performs action augmentations within the finite symbolic space to produce a Geo-DSL sequence. This approach effectively constrains the synthesis and augmentation process, ensuring validity and controllability by generating outside the infinite domains of natural language and image spaces (Section 3.2). Second, a conversion engine maps the generated Geo-DSL sequences back into natural language descriptions and visual image representations. This process synthesizes high-quality images and valid text while avoiding intermodal information overlap. Third, to get Q&A Pairs with reasoning paths, we utilize expert LLMs to conduct backwards search to identify the geometric unknowns to be solved and generate the CoT in a forward manner (Section 3.4). The search process primarily ensures the diversity of the Q&A pairs, while the forward verification confirms the correctness of the CoT and the final answer. The overall framework of NeSyGeo is illustrated in Figure 3.
3.1 Symbolic Definition
Existing symbolic languages for plane geometry have certain limitations. The definitions in AlphaGeometry [23] are tailored for proof-based problems, thus lacking definitions related to specific numerical values. InterGPS [18] defines a predicate as a geometric shape entity, geometric relation, or arithmetic function, constructing 91 predicates. However, this approach overly fragments shape, attribute, and relation definitions into independent statements, often requiring multiple statements to specify a single element in the figure. This significantly increases the complexity of a conversion engine’s identification of elements within the symbolic space and further conversion of them.
We propose Geo-DSL, a concise and comprehensive symbolic language for plane geometry to address these limitations. It employs an entity-relation-constraint framework, using well-defined rules to uniquely define 13 types of points, 7 types of lines, 3 types of angles, and 14 types of shapes, covering all plane geometry elements and incorporating numerical attributes like lengths and angles for precise specifications. Table 1 provides partial examples of symbolic definitions. Geo-DSL offers two key advantages. First, its comprehensive coverage includes all geometric elements and their numerical properties, enabling accurate and complete descriptions. Second, its simplicity allows a single statement to specify an element uniquely, promoting the incremental integration of the symbolic action space and the sequential parsing of statements by the conversion engine. This combination of completeness and simplicity makes Geo-DSL an efficient and powerful geometric representation and processing solution.
3.2 Symbolic Sequence Generation
To generate a Geo-DSL sequence, we introduce a step-action augmenter that iteratively synthesizes a sequence of statements, as detailed in Algorithm 1. Based on dataset preferences, we first configure the step count $N$, weight matrices $I$ and $A$ for selecting elements and actions with respective probabilities, and ranges $[l_{\text{min}},l_{\text{max}}]$ for lengths and $[\theta_{\text{min}},\theta_{\text{max}}]$ for angles. Then, the augmenter iteratively generates symbolic statements over $N$ steps. For each step, we randomly sample parameters $x,y,z$ and $\alpha$, select an element $v_{j}\in f_{v}$ using weights from $I$, and choose an action $a_{k}$ based on weights from $A$ (see Appendix I for action details). The new statement $s_{\text{new}}$ is then incorporated into the sequence $f_{s}$. Leveraging Geo-DSL’s symbolic definitions and well-defined actions, this approach ensures the validity and accuracy of each statement. Meanwhile, randomized elements, diverse action selections, and customizable hyperparameter preferences promote diversity.
Algorithm 1 The overall framework of the symbolic sequence generation process
1:Input: Step count $N$, Action weight matrix $A$, Element selection weight matrix $I$, Line length range $[l_{min},l_{max}]$, Angle range $[\theta_{min},\theta_{max}]$$\triangleright$ Set customizable hyperparameter
2:Output: Generated Geo-DSL statement sequence $f_{s}$.
3:Initialize $f_{s}$=Initialize( ) . $\triangleright$ Initialize the sequence $f_{s}$ with the first statement
4:Initialize symbolic state space elements $f_{v}=\text{Initialize}(f_{s})$. $\triangleright$ Initialize state based on $f_{s}$
5:for $i=1$ to $N$ do
6:     Randomly sample $x,y,z$ from $[l_{min},l_{max}]$.
7:     Randomly sample $\alpha$ from $[\theta_{min},\theta_{max}]$.
8:     Element $v_{j}$=Selected_elements($f_{v}$,$I$). $\triangleright$ Select element $v_{j}$ from $f_{v}$ randomly.
9:     Action $a_{k}$=Selected_action($v_{j}$,$A$). $\triangleright$ Select action $a_{k}$ based on the type of $v_{j}$ randomly
10:     $s_{new}$=Generate_DSL($a_{k}$, x, y, z, $\alpha$). $\triangleright$ Generate new DSL statement
11:     $f_{s}$=Update($f_{s}$,$s_{new}$) $\triangleright$ Add the new statement to the sequence
12:     $f_{v}$=Update($f_{v}$,$s_{new}$). $\triangleright$ Update state space elements
13:end for
14:return $f_{s}$.
3.3 Informalization
Following the generation of sequences within the formal symbolic space, it is necessary to map them back to the text and image spaces for further processing and visualization. Our approach generates high-quality images and rigorous natural language, allocating information between them to compel MLLMs to leverage visual data effectively.
Figure 4: Reverse search and forward validation with expert LLMs
Visual image. For the visual space, our visualization engine parses each Geo-DSL statement as a geometric element to generate high-quality images using Matplotlib. The rigorous symbolic language space ensures the determinacy and uniqueness of the conversion process, thereby enabling precise image generation. Additionally, the engine produces images with detailed annotations absent from the text, requiring models to leverage images during problem-solving, enhancing their visual perception and ability to extract image-based information.
Natural language. We adopt a template-based transformation approach, parsing and mapping each symbolic statement to multiple predefined natural language templates. This produces a text-full version containing all details and captions of the image and a text-lite version retaining only essential conditions, not included in the image annotations, as the final form in our datasets. The text-lite version avoids intermodal information redundancy, while diverse templates ensure the validity and diversity of the synthesized condition text.
3.4 CoT Generation
To generate Q&A pairs with CoT reasoning, we utilize the strong reasoning abilities of expert LLMs to ensure diverse and reliable output. Using the text-full version generated as described in Section 3.3 as input, we develop a two-step process comprising reverse search and forward validation as shown in Figure 4. Prompts are detailed in Appendix G.
Reverse search. We use DeepSeek R1 [6] for reverse search, starting from the given conditions, iteratively exploring and deriving conclusions step by step, and ultimately producing the final conclusions at the end of the reasoning chain along with their corresponding answers. This reverse approach starts from known information and recursively builds a reasoning chain, reducing Q&A generation complexity and hallucinations. R1’s strong reasoning and exploration capabilities yield diverse conclusions, enriching the variety of Q&A pairs.
Forward validation. To ensure the correctness of Q&A pairs and generate step-by-step CoT reasoning, we re-input the questions and text-full conditions into DeepSeek V3, requesting both the reasoning process with final answer, and cross-validate these answers against those from R1 to include only consistent pairs in our final dataset. This process guarantees answer correctness and yields diverse, valid CoT without an extensive search space.
4 Experiments
We present a series of experiments to investigate the following four research questions:
Efficacy – To what extent does training on our synthesized dataset in both RL and SFT improve the geometric reasoning performances of several MLLMs?
Efficiency – Is the data generated by NeSyGeo more effective in yielding models with better performance compared to using data from existing automatic synthesis frameworks?
Diversity – Does the NeSyGeo framework effectively ensure diversity across both the text and image spaces of the generated synthetic geometric dataset?
Visual Effectiveness – Can our datasets compel models to effectively utilize visual information for enhanced understanding by appropriately distributing information between modalities?
4.1 Experimental Setup
Dataset. To synthesize a diverse dataset, we set hyperparameters for generating images by configuring the step count, length range, and angle range. The step count $N$ ranges from one to four. The line length is defined within the basic range $[l_{\text{min}},l_{\text{max}}]=[1,5]$, scalable by any multiple of 2 (e.g., $[4,20]$). The angle is constrained to multiples of 15° within $[15^{\circ},165^{\circ}]$, with increased weights assigned to special angles. We generate various types of weight matrices $A$ and $I$ by adjusting their corresponding values. This process yields a NeSyGeo-CoT dataset with 30k Q&A pairs and a NeSyGeo-Caption dataset with 70k Q&A pairs. Additional dataset statistics are provided in Appendix B.
Evaluation. Our evaluation is conducted on several benchmarks: the Test set of GeoQA[4], the Test_MINI set of MathVision[26], and the MathVerse[36]. For the MathVerse benchmark, we select the Vision Only, Vision Dominant, and Vision Intensive sets to better assess the visual perception and logical reasoning capabilities of MLLMs. We extract in-domain metrics from other datasets, including angle, area, length, and Plane Geometry, to effectively evaluate the models’ capabilities in geometric reasoning problems, in addition to the GeoQA dataset, which focuses entirely on plane geometry. For GeoQA, we employed hard-coded extraction for comparison, while other evaluations are assessed using the automated VLMEvalKit framework[8]. Appendix C provides additional experimental details and evaluation results.
4.2 Empirical Results
Efficacy: Training multiple MLLMs with NeSyGeo via both SFT and RL significantly enhances geometric problem-solving performances.
We first sample 4k samples from our NeSyGeo-CoT dataset and apply the Group Relative Policy Optimization (GRPO) algorithm to train two epochs with Deepseek R1’s format and answer rewards. The training code framework is based on VLM-R1 [22]. As shown in Tables 2, models achieve the best performance among the baselines after training. InternVL2.5-4B significantly improved in the angle knowledge domain with gains of 8.4 (MathVerse), 7.3 (GeoQA), and 5.3 (MathVision). Qwen2.5-VL-3B achieved a +15.8 performance boost in the area domain of MathVision. Notably, across all evaluated metrics, the InternVL2.5-4B model trained on the NeSyGeo dataset achieves performance on par with or superior to its 8B counterpart.
We also conducted SFT experiments, initially training on our NeSyGeo-Caption dataset to enhance the models’ perception of geometric images, followed by training on the NeSyGeo-CoT dataset to improve reasoning capabilities. The experiments were conducted on LLaMA-Factory [40] framework. Evaluation results on MathVerse (Vision Intensive) and GeoQA are presented in Table 3. The trained model demonstrates performance improvements over the base model on most metrics.
Efficiency: Under the same data budget, the generated data from our framework is better than that from popular automatic generation frameworks.
We randomly sampled 4k samples from MAVIS [37] and R-CoT [7], which are automatic frameworks in geometry problem generation. To ensure a fair comparison, we maintained consistent settings.
Table 2: RL performance comparison: Models trained with only 4k samples of NeSyGeo-CoT show performance gains over the base models, with the InternVL2.5-4B model exceeding the 8B variant in geometry problem-solving.
GeoQA MathVision MathVerse
Model Angle Area Length Angle Area Length Plane Geometry
Qwen2.5-VL-3B 53.3 26.3 26.3 21.1 31.3 20.9 37.0 32.5
Qwen2.5-VL-3B+NeSyGeo 55.7 (+2.4) 26.3 (+0.0) 42.1 (+15.8) 26.3 (+5.2) 32.6 (+1.3) 23.5 (+2.6) 37.2 (+0.2) 35.5 (+3.0)
InternVL2.5-4B 61.9 36.8 31.6 26.3 31.5 22.7 31.9 30.7
InternVL2.5-4B+MAVIS 63.5 (+1.6) 31.6 (-5.2) 26.3 (-5.3) 31.6 (+5.3) 37.1 (+5.6) 20.9 (-1.8) 35.3 (+3.4) 33.7 (+3.0)
InternVL2.5-4B+R-CoT 63.3 (+1.4) 31.6 (-5.2) 31.6 (+0.0) 21.1 (-5.2) 31.2 (-0.3) 18.3 (-4.4) 34.3 (+2.4) 28.7 (-2.0)
InternVL2.5-4B+NeSyGeo 69.2 (+7.3) 42.1 (+5.3) 36.8 (+5.2) 26.3 (+0.0) 39.9 (+8.4) 24.9 (+2.2) 36.1 (+4.2) 36.7 (+6.0)
InternVL2.5-8B 66.2 36.8 36.8 21.1 36.9 23.1 34.8 36.6
Table 3: SFT performance comparison: The trained model demonstrates performance improvements over the base model on most metrics.
GeoQA Vision Intensive
Model Angle Area Length Plane Geometry
Qwen2.5-VL-7B 69.4 43.0 27.5 46.2 44.1
Qwen2.5-VL-7B+NeSyGeo 71.8 (+2.4) 46.1 (+3.1) 23.1 (-4.4) 49.5 (+3.3) 46.7 (+2.6)
LLaVA-Next-7B 22.6 28.5 6.6 16.5 20.4
LLaVA-Next-7B+NeSyGeo 26.1 (+3.5) 30.6 (+2.1) 7.7 (+1.1) 19.2 (+2.7) 22.9 (+2.5)
As illustrated in Figure 5 and Table 2, while all datasets help the model improve over the baseline, our dataset outperformed others in most metrics, validating its superior performance.
Diversity: The dataset synthesized by our NeSyGeo framework exhibits high diversity in both text and visual features.
A critical challenge for automatic data synthesis methods is whether the dataset is sufficiently diverse to avoid quality degradation due to potential overfitting risks from inherent domain constraints. We employ t-SNE [24] dimensionality reduction for mapping in text space to evaluate the diversity across different methods. This analysis allows us to assess the diversity of the textual descriptions themselves. Given that in geometric problems, the text conditions depict specific visual elements, and the diversity observed in the text space also serves as a valuable indicator of the diversity in the corresponding visual diagrams. To ensure a fair comparison, we remove all prompts related to guiding large models, retaining only condition and question texts, and randomly sample 5k texts from each dataset. The results are illustrated in Figure 6. Our method and G-LLaVA [9] exhibit uniformly distributed features in the space, indicating low data overlap and high diversity. In contrast, R-CoT and MAVIS display varying degrees of clustered distribution, indicating more feature-similar samples. To directly assess the diversity of the visual features, we also performed t-SNE on image features extracted by ResNet, with detailed experimental results presented in Appendix C.
Figure 5: Efficiency comparison of our NeSyGeo-CoT dataset versus other mainstream automated synthesis datasets. The models are trained using RL methods with InternVL2.5-4B.
Figure 6: T-SNE of the text features of different automatic frameworks. The G-LLaVA method augments the text space on the manually annotated GeoQA dataset. Thus, its text diversity can approximate that of real data more closely. Similar to G-LLaVA, our method exhibits a uniform distribution in the space, demonstrating superior diversity.
Table 4: Comparison between NeSyGeo-CoT and text-redundant datasets with equivalent data budgets. Models are trained via RL on the InternVL2.5-4B. The highest value for each metric is underlined. The results show that our dataset improves models’ visual perception and logical reasoning capabilities.
Text Dominant Vision Only
Dataset Angle Area Length Plane Geometry Angle Area Length Plane Geometry
Base 47.1 27.5 43.4 44.1 24.4 20.9 29.1 27.3
NeSyGeo 49.2 28.6 44.0 45.5 29.0 27.5 34.1 31.8
NeSyGeo+RED 52.8 27.5 44.5 45.1 27.5 25.3 33.0 30.2
R-CoT 51.8 24.2 45.0 46.5 25.9 23.1 31.3 29.2
Visual Perception: Models trained on NeSyGeo data shows modest gains over information-redundant datasets when textual shortcuts exist, yet achieves substantial improvements when image understanding is necessary. This indicates that ours enhances not only logical reasoning but also image perception and utilization of the model.
A key question is whether reducing redundancy and forcing models to extract visual information improves their geometric reasoning capabilities. We evaluate models on the MathVerse(Text Dominant), which provides redundant text descriptions and implicit properties enabling reasoning without images, and the MathVerse(Vision Only) version, where all information is embedded entirely within the images. For this comparison, we selected two text-redundant datasets: NeSyGeo+RED, the original NeSyGeo-CoT dataset supplemented with textual equivalents of its image annotations, and the R-CoT dataset. Results presented in Table 4 show that our model outperforms the baseline on Text-Dominant but lags behind other datasets in some metrics. On Vision-Only, our model surpasses them across all metrics, demonstrating enhanced geometric reasoning and visual perception.
5 Conclusion
This paper introduces NeSyGeo, a neurosymbolic framework for automatically synthesizing multimodal geometric datasets. Our approach transforms the generation process into a controllable symbolic space using Geo-DSL, maps the symbolic representation back to image and natural language spaces via a conversion engine, and then utilizes LLMs for backwards search and forward solving to produce Q&A pairs. Using this framework, we construct the NeSyGeo-CoT and NeSyGeo-Caption datasets, totalling 100k samples. We also propose NeSyGeo-Test, a comprehensive benchmark for evaluating MLLMs’ geometric reasoning capabilities. Our datasets significantly and consistently improve the reasoning abilities of multiple MLLMs through both SFT and RL.
Future Work: We intend to extend NeSyGeo to other multimodal domains, such as analytical geometry and visual question answering. This extensibility will be achieved by defining new domain-specific languages, corresponding synthesis rules within the symbolic space, and tailored conversion engines. Furthermore, we plan to develop an automated symbolic solver capable of conducting search and validation directly within the symbolic space. This would remove reliance on LLMs, potentially reducing generation costs and ensuring complete correctness of the datasets.
Appendix A Comparison with Specific Examples of Popular Geometry Datasets
To facilitate comparison of dataset characteristics synthesized by our method and other popular approaches, we showcase a randomly selected example from NeSyGeo-CoT alongside each of the different approaches in Figure 7. Geometry-3K is a manually synthesized dataset, while the remaining approaches employ automatic generation techniques. To ensure a fair comparison, we standardize the text format by removing model-guiding prompts and appending options when present. Furthermore, we annotate each sample with image pixels and CoT word counts.
Compared to other datasets, our dataset features clear, human-aesthetically pleasing images, high-quality step-by-step reasoning chains, symbolic form meta-information enabling subsequent image augmentation and mutation, and well-distributed conditional information between images and text. Additional examples of our NeSyGeo-CoT dataset can be found in the Appendix 11.
Figure 7: Comparison NeSyGeo-CoT dataset with other Popular Geometry Datasets. Geometry-3K is a manually synthesized dataset, while the remaining approaches employ automatic generation techniques. Our dataset features clear, human-aesthetically pleasing images, high-quality step-by-step reasoning chains, symbolic form meta-information enabling subsequent image augmentation and mutation, and well-distributed conditional information between images and text.
Appendix B Statistics of NeSyGeo-Caption and NeSyGeo-CoT
Detailed numerical statistics and element distribution for the NeSyGeo-Caption and NeSyGeo-CoT datasets are presented in Table 6 and  6. For element distribution statistics, we randomly sampled 1.8k Geo-DSL sequences corresponding to images from each dataset, counting the frequency of different geometric elements. To facilitate interpretation, these elements are converted into corresponding natural language descriptions.
Table 5: Statistics of NeSyGeo-Caption
Statistic Number
Total Counts
Total number of images 70k
Total number of captions 70k
DSL Statement Percentage
One Statement 5.4%
Two Statements 25.8%
Three Statements 34.7%
Four Statements 23.4%
Five Statements 10.7%
Length of Captions
Maximum length (words) 220
Minimum length (words) 34
Average length (words) 73.3
Average length (characters) 385.4
Image Dimensions
Average dimensions (pixels) 723.1 $\times$ 724.0
Table 6: Statistics of NeSyGeo-CoT
Statistic Number
Total Counts
Total number of images 15.3k
Total number of Q&A pairs 30.1k
Question Statistics
Length-based type 54.6%
Area-based type 34.4%
Angle-based type 11.1%
Average length (words) 26.9
Average length (characters) 140.6
CoT Statistics
Below four steps 35.4%
Four steps or above 64.5%
Average length (words) 91.8
Average length (characters) 365.9
Image Dimensions
Average dimensions (pixels) 731.0 $\times$ 727.4
Figure 8: Frequency of different geometric elements. To facilitate interpretation, these elements are converted into corresponding natural language descriptions.
Appendix C Additional Experimental Details and Results
We utilized the VLM-R1 [22] framework for RL experiments, conducted on 6 vGPU-32 GB. We set epochs to 2, num generations to 6, batchsize to 1. To enhance the visual perception capabilities of MLLMs, parameters of the language model and vision modules are set to be trainable.
Table 7: Detailed RL experiments evaluation on MathVerse. Here, ‘AGL’, ‘ARA’, ‘LTH’, and ‘PG’ denote angle, area, length, and plane geometry, respectively.
Vision Intensive Vision Dominant Vision Only
Model AGL ARA LTH PG AGL ARA LTH PG AGL ARA LTH PG
Qwen2.5-VL-3B 31.6 22.0 34.6 33.3 31.6 17.6 40.7 31.4 30.6 23.1 35.7 32.7
InternVL2.5-4B 36.8 22.0 33.0 31.8 33.2 25.3 33.7 32.9 24.4 20.9 29.1 27.3
InternVL2.5-8B 44.0 23.1 36.3 41.8 40.4 20.9 36.8 37.3 26.4 25.3 31.3 30.6
Qwen2.5-VL-3B+RL 33.7 23.1 36.8 34.7 32.1 20.9 37.4 36.6 32.1 26.4 37.4 35.1
InternVL2.5-4B+RL 45.6 20.9 37.9 40.2 45.1 26.4 36.2 38.2 29.0 27.5 34.1 31.8
For SFT experiments, we employed the LLaMA-Factory [40] framework on 2 A800 GPUs with LoRA. We set the learning rate of $1\times 10^{-5}$, LoRA rank of 64, and use Adam optimization. Training on NeSyGeo-Caption used 1 epoch, while NeSyGeo-CoT used 2 epochs.
Table 7 presents the detailed performance of models trained on various automatically synthesized datasets across the MathVerse benchmark. Models trained using our dataset demonstrate superior performance on most metrics compared to others, exhibiting substantial performance gains relative to the base model.As shown in Figure 9. We also evaluated model performance as RL training steps increased when using NeSyGeo-CoT. Most metrics improved with more training steps, demonstrating the robustness and effectiveness of our datasets.
To investigate visual diversity directly across datasets, we randomly sampled 1k images from each, extracted features using ResNet, and visualized them via t-SNE. As illustrated in Figure 10, G-LLaVA—having augmented only the textual components of its base dataset—displays a distinctly non-uniform distribution in the image feature space. Conversely, our method exhibits uniform feature distributions, underscoring the substantial visual diversity of images generated by our approach.
Figure 9: Model performance on Mathverse as the RL training steps increase. With InternVL2.5-4B as our base model, most metrics exhibit progressive improvement throughout training, demonstrating the robustness and effectiveness of our datasets.
Figure 10: T-SNE of the image features of different automatic frameworks. Our datasets exhibit uniform feature distributions, underscoring the substantial visual diversity of images generated by the NeSyGeo framework.
Appendix D More Examples of NeSyGeo-CoT Dataset
Figure 11: Examples of the NeSyGeo-CoT dataset. Each sample comprises a symbolic image definition based on our Geo-DSL language, a high-quality annotated image, a concise text caption, diverse Q&A pairs, and a detailed reasoning process step-by-step.
We present more examples from the NeSyGeo-CoT dataset in Figure 11. Our bidirectional conversion engine can generate high-quality visual images from a symbolic form based on our Geo-DSL language. To further enhance image diversity, the engine introduces variability by randomly selecting values for the unit length and applying random rotations to the generated diagrams during creation. While other visual attributes, such as element and background colours, could also be randomized, they were set to default values in our current synthesis process. Our symbolic language helps identify parts of the image, and our conversion process ensures the images are geometrically correct.
Due to LLMs’ powerful search and reasoning capabilities, we obtain diverse Q&A pairs concerning properties like lengths, angles, or areas alongside high-quality CoT step-by-step. This process, involving backwards search across the geometric space defined by the symbolic form and forward validation, ensures the correctness of the numerical answers, thereby enriching textual diversity.
To support evaluation and training paradigms such as curriculum learning, we annotate each sample with a difficulty level. Given that geometric reasoning tasks primarily require models’ image perception and logical reasoning capabilities, we scientifically define the difficulty level as
[…]
where perception difficulty is the number of Geo-DSL statements, and reasoning difficulty is the number of reasoning steps.
Each synthesized sample includes detailed meta-information stored as a symbolic form based on our Geo-DSL language. This symbolic form accurately describes the geometric setup and offers promising directions for future research. For instance, valid geometric configurations could be generated by augmenting or mutating existing symbolic forms within constrained parametric bounds.
Appendix E Details of NeSyGeo-Test Benchmark.
Our NeSyGeo-Test benchmark comprises 2668 Q&A pairs. Consistent with the training set, numerical annotations are embedded in the image space, with only essential conditions and questions provided in the text. The type of numerical quantity categorizes the dataset sought: Angle (658 pairs), Shape (730 pairs), and Length (1280 pairs). Shape type includes shape area and perimeter, while length type includes edge and arc lengths. Based on the difficulty level in  D, problem difficulty is divided into three levels: Easy (1537 pairs), Medium (908 pairs), and Hard (223 pairs). Evaluation results on current mainstream open-source and closed-source MLLMs are shown in Table 8.
Table 8: NeSyGeo-Test Benchmark on several mainstream MLLMs. The highest accuracy for open-source and closed-source MLLMs is marked in red and blue, respectively.
Task Type Question Difficulty
Model Param Angle Shape Length Easy Medium Hard Total
Open-source MLLMs
Qwen2.5-VL-3B-Instruct 3B 44.1 31.8 27.8 30.5 29.1 31.9 30.9
Qwen2.5-VL-7B-Instruct 7B 38.3 30.7 32.2 36.8 27.5 32.3 43.3
InternVL2.5-4B 4B 53.8 53.1 53.5 62.9 49.4 32.0 53.4
InternVL2.5-8B 8B 55.9 56.2 55.4 64.5 49.4 43.3 55.8
LLaVA-NeXT-7B 7B 23.7 15.5 15.4 18.6 14.7 13.3 6.4
LLaVA-NeXT-13B 13B 15.7 15.6 16.0 15.8 16.0 15.2 15.8
LLaVA-NeXT-34B 34B 23.7 21.0 18.5 19.1 21.5 19.2 19.9
Closed-source MLLMs
InternVL3-latest 81.7 65.2 68.3 77.8 62.0 56.7 68.7
GPT-4o-mini 58.7 62.1 55.7 63.0 54.0 41.7 58.2
Claude-3.5-Sonnet-latest 68.8 78.0 71.0 77.8 73.2 56.5 74.5
Qwen-VL-plus 38.5 29.6 31.7 36.6 27.2 29.6 32.8
Gemini-2.0-Flash 36.8 60.4 67.7 54.3 63.8 61.0 58.1
Appendix F Limitations
  • Limited Training Paradigm: Our current evaluation of the NeSyGeo dataset relies on a simple training paradigm to assess its efficacy for automated data generation. This approach lacks advanced training strategies, such as CLIP alignment or curriculum learning, which restricts the development of specialized models optimized for geometric reasoning tasks.
  • Restricted Domain Scope: The NeSyGeo framework is currently tailored to plane geometry, limiting its generalizability to other domains. However, we believe that for multimodal datasets in other domains, we can similarly achieve synthesis by defining symbolic statements, shifting the synthesis process to a controllable symbolic space, and constructing a symbolic-to-image engine, which we plan to explore in our future work.
  • Dependency on External APIs: The construction of Q&A pairs in this study partly relies on the reasoning capabilities of LLMs. This dependency increases generation costs and introduces potential inconsistencies. We aim to develop an automated solver that conducts search and validation directly within the symbolic space, thereby removing reliance on LLMs. This could further reduce costs and ensure complete rigor.
Appendix G Details of Prompts in Reverse Search and Forward Validation
In our automatic synthesis framework, we employ DeepSeek R1 as the expert LLM for reverse search and DeepSeek V3 for forward validation. The specific prompts utilized are detailed in Figures 12 and 13, respectively. Note that the blue text in these prompts is substituted with actual content.
Appendix H Detailed Definition of Geo-DSL
Geo-DSL adopts an entity-relation-constraint framework to define geometric elements in plane geometry, encompassing 13 types of points, 7 types of lines, 3 types of angles, and 14 types of shapes. Representative examples of symbolic statements and their corresponding natural language descriptions are illustrated in Figures 18, 15, 16, and 17. With a single statement, Geo-DSL uniquely specifies spatial elements, ensuring the accuracy of geometric synthesis while significantly facilitating parsing and transformation by our conversion engine. This language achieves comprehensive coverage of plane geometry, including numerical attributes such as lengths and angle measures, enabling precise and complete geometric representations. By streamlining definitions into concise statements, Geo-DSL reduces the complexity of symbolic processing, enhances the efficiency of the conversion engine, and supports seamless integration with neural synthesis pipelines. These advantages make Geo-DSL a robust and versatile solution for generating high-quality multimodal geometric reasoning data.
Appendix I Detailed Actions in Symbolic Spaces
As illustrated in Figure 14, we enumerate all statements within the action space defined by our Geo-DSL. The content within square brackets denotes annotations for each statement.
As outlined in Algorithm 1, for each step, we first generate three lengths, $x$, $y$, and $z$, along with an angle $\alpha$, sampled from predefined ranges. Subsequently, based on the weight matrices $A$ and $I$, we determine the specific statement to be selected. The chosen statement, paired with its corresponding numerical values, is then appended to the Geo-DSL sequence. For different types of actions, we provide a concrete example for each, highlighted with a gray background to indicate the available action space when the respective element is selected.
Figure 12: Prompt for Deepseek R1 in reverse search.
Figure 13: Prompt for Deepseek V3 in forward validation.
Figure 14: Detailed Actions in Symbolic Spaces. Actions can be categorized into four parts based on the type of selected geometric element: line-based, point-based, shape-based, and angle-based.
Figure 15: Geo-DSL definitions of line.
Figure 16: Geo-DSL definitions of angle.
Figure 17: Geo-DSL definitions of shape.
Figure 18: Geo-DSL definitions of point.

### v2

@@
 Abstract
-Obtaining large-scale, high-quality data with reasoning paths is crucial for improving the geometric reasoning capabilities of multi-modal large language models (MLLMs).
+Obtaining large-scale, high-quality reasoning data is crucial for improving the geometric reasoning capabilities of multi-modal large language models (MLLMs).
 However, existing data generation methods, whether based on predefined templates or constrained symbolic provers, inevitably face diversity and numerical generalization limitations.
 To address these limitations, we propose NeSyGeo, a novel neuro-symbolic framework for generating geometric reasoning data.
-First, we propose a domain-specific language grounded in the entity–relation–constraint paradigm to comprehensively represent all components of plane geometry, along with generative actions defined within this symbolic space.
-We then design a symbolic–visual–text pipeline that synthesizes symbolic sequences, maps them to corresponding visual and textual representations, and generates diverse question–answer (Q&A) pairs using large language models (LLMs).
-To the best of our knowledge, we are the first to propose a neuro-symbolic approach in generating multimodal reasoning data.
+First, we propose a domain-specific language grounded in the entity–attributes-relations paradigm to comprehensively represent all components of plane geometry, along with generative actions defined within this symbolic space.
+We then design a symbolic–visual–text pipeline that synthesizes symbolic sequences, maps them to visual and textual representations and generates reasoning path with reverse search and forward validation.
 Based on this framework, we construct NeSyGeo-CoT and NeSyGeo-Caption datasets, containing 100k samples, and release a new benchmark NeSyGeo-Test for evaluating geometric reasoning abilities in MLLMs.
@@
 1 Introduction
-Improving the visual reasoning capabilities of MLLMs has garnered significant attention recently  [17, 2, 1, 14, 11, 33, 25, 28, 15], with models like InternVL [5] and the QwenVL series [27, 3] demonstrating significant enhancements in visual-semantic comprehension through their multimodal capabilities.
-Among various visual reasoning tasks, geometric mathematical reasoning is crucial for evaluating the reasoning performance of MLLMs [31, 30], as it requires a deep integration of spatial perception, symbolic understanding, and logical deduction.
-To enhance such reasoning abilities, existing approaches [37, 38, 39] primarily rely on fine-tuning base models using reinforcement learning (RL) or supervised fine-tuning (SFT) on specialized geometric reasoning datasets.
+Improving the visual reasoning capabilities of MLLMs has garnered significant attention recently  (Liu et al., 2023; Alayrac et al., 2022; Achiam et al., 2023; Li et al., 2021; Jiang et al., 2024; Zhang et al., 2025a; Wang et al., 2024a; Wu et al., 2024; Liang et al., 2024), with models like InternVL (Chen et al., 2024) and the QwenVL series (Wang et al., 2024c; Bai et al., 2025) demonstrating significant enhancements in visual-semantic comprehension through their multimodal capabilities.
+Among various visual reasoning tasks, geometric mathematical reasoning is crucial for evaluating the reasoning performance of MLLMs (Yan et al., 2025; 2024), as it requires a deep integration of spatial perception, symbolic understanding, and logical deduction.
+To enhance such reasoning abilities, existing approaches (Zhang et al., 2024d; 2025b; Zhao et al., 2025) primarily rely on fine-tuning base models using reinforcement learning (RL) or supervised fine-tuning (SFT) on specialized geometric reasoning datasets.
 These methods depend heavily on the availability of large-scale, high-quality geometric reasoning data, which is often costly and time-consuming to construct manually.
 Therefore, automatic data generation for geometric reasoning has emerged as a promising and actively explored direction, aiming to alleviate data scarcity and further improve the reasoning abilities of MLLMs.
-Existing approaches for generating datasets in geometric tasks can be broadly classified into four categories. Text augmentation methods like G-LLaVA [9] primarily mutate the conditions of
 Figure 1: Performance comparison of different MLLMs and LLMs with and without image input in several geometry datasets.
 The minimal or negligible drops observed upon image removal in GeoQA and R-CoT raise concerns regarding the utilization of visual information for geometric reasoning.
-existing datasets through equivalent condition transformation and numerical scaling.
+Existing approaches for generating datasets in geometric tasks can be broadly classified into four categories.
+Text augmentation methods like G-LLaVA (Gao et al., 2025) primarily mutate the conditions of existing datasets through equivalent condition transformation and numerical scaling.
 However, this approach fails to address the scalability of image generation.
-Template-based methods [7, 37, 12], use predefined geometric templates with fixed topologies, simplifying synthesis but constraining diversity by reducing the geometric space to limited combinations.
-Solver-based methods [10, 32] inspired by symbolic prover AlphaGeometry [23], leverage formal languages for synthesis but lack metric details (e.g., angles, lengths, areas), restricting multimodal data to descriptive annotations and limiting numerical reasoning applications.
-Tool-based methods attempt to generate codes for tools like GeoGebra or MATLAB via LLMs.
-However, even advanced models struggle to ensure correctness with ambiguous natural language instructions and complex geometric spaces.
-In summary, existing methods grapple with issues of image scalability, limited geometric diversity, a lack of precise numerical information, and challenges in ensuring the reliability of generated content.
-Beyond the challenges in data synthesis methodologies, current geometric reasoning datasets present several limitations that impede the advancement of MLLMs.
-A primary limitation stems from the often inadequate quality and low resolution of the provided images.
-Such inputs frequently fall below the optimal requirements of visual encoders [20, 16, 17], hindering the extraction of crucial fine-grained visual features and discriminative information essential for robust multimodal reasoning.
-Furthermore, our analysis reveals notable information redundancy between the textual and visual modalities in many current datasets.
+Template-based methods (Deng et al., 2024; Zhang et al., 2024d; Kazemi et al., 2024), use predefined geometric templates with fixed topologies, simplifying synthesis but constraining diversity by reducing the geometric space to limited combinations.
+Solver-based methods (Huang et al., 2025; Zhang et al., 2024a) inspired by symbolic prover AlphaGeometry (Trinh et al., 2024), leverage formal languages for synthesis but lack metric details (e.g., angles, lengths, areas), restricting multimodal data to descriptive annotations and limiting numerical reasoning applications.In summary, existing methods grapple with issues of image scalability, limited geometric diversity, a lack of precise numerical information, and challenges in ensuring the reliability of generated content.
+Moreover, current geometric reasoning datasets present three critical limitations in cross-modal reasoning:1) Information Redundancy: Repetitive textual conditions provide a “shortcuts” that allows models to bypass visual reasoning.
+2) Semantic Gap Between Modalities: Abstract geometric properties stated in the text (such as equality) are not explicitly grounded from the visual features.
+3) Poor Image Quality: Low-resolution and inadequate images hinder the extraction of crucial visual information.
 As shown in Figure 1, our comparative experiments demonstrate minimal or negligible accuracy drops upon image removal.
-This finding emphasizes the urgent need for a dataset that effectively separates textual and visual information and provides high-quality images to promote MLLMs’ visual perception and logical reasoning performance.
+Such discrepancies hinder the model’s ability to make use of visual perception and learn robust cross-modal reasoning .
 To address these challenges, we propose NeSyGeo, a neuro-symbolic framework for synthesising high-quality multimodal geometric reasoning datasets.
-NeSyGeo integrates three components: 1) A formal geometric symbolic space defined by a domain-specific language (DSL), capturing primitive entities (points, lines, circles), topological relations (parallelism, incidence, perpendicularity), and metric constraints (angles, lengths), enabling diverse geometric configurations via systematic sampling within constrained parametric bounds.
-2) A bidirectional conversion engine that transforms symbolic constructs into decoupled modalities, producing annotated vector graphics paired with concise textual axioms.
-3) A causal Q&A pairs and theorem-grounded Chain-of-Thought (CoT) sequences generator that effectively merges neural reasoning with symbolic verification.
-To our knowledge, we are the first to develop a neuro-symbolic framework for producing multimodal reasoning data.
-Our framework enhances the diversity and validity of generated geometric reasoning data while effectively mitigating information redundancy and the underutilization of visual signals during training.
-Specifically, the comprehensive Geo-DSL and its expansive symbolic synthesis action-space promote diverse and well-grounded image generation.
-Meanwhile, our CoT sequence generator, powered by LLMs’ strong reasoning and language capabilities, conducts a backwards search across the geometric space to construct Q&A pairs, thereby enriching textual diversity.
-The unique identification of geometric elements via our symbolic language and dedicated conversion engine ensures visual validity.
-In parallel, a bidirectional cross-validation process using expert LLMs ensures textual validity.
-By strategically distributing complementary information across image and text modalities, our approach
+NeSyGeo integrates three components: 1) A formal geometric symbolic space defined by a domain-specific language (DSL), capturing primitive entities (points, lines, circles), metric attributes (angles, lengths) and topological relations (parallelism, incidence, perpendicularity), enabling diverse geometric configurations via systematic sampling within constrained parametric bounds.
+2) Bidirectional conversion engines that transform symbolic constructs into decoupled modalities, producing annotated vector graphics paired with concise textual axioms.
+3) Validated Question-Answer (Q&A) pairs and theorem-grounded CoT generators that effectively merges neural reasoning with verification.
+Our framework enhances data quality by generating a wide spectrum of well-grounded images and correspondingly complex textual descriptions, while ensuring the validity and fidelity of both modalities.
+Furthermore, by carefully annotating semantic relationships within the image and avoiding redundancy between textual and visual conditions, it compels MLLMs to actively engage with visual information, thereby preventing modality neglect and significantly improving their visual grounding capabilities during training.
+[…]
 Figure 2: Comparison of dataset characteristics synthesized by our method and other popular synthesis approaches.
-“High Resolution” denotes average image pixels exceeding 336$\times$336.
+“High Resolution” denotes average image pixels exceeding $\times336$.
 “Symbolic Form” refers to the symbolic meta-information associated with the image.
@@
 “Visual Understanding” represents the mitigation of image-text redundancy for stronger visual grounding in reasoning.
-More specific examples of different methods are in Appendix A.
-encourages MLLMs to actively engage with visual information when solving problems, enhancing their abilities to perceive visually and effectively utilize images.
-Leveraging the NeSyGeo pipeline, we construct two training datasets, NeSyGeo-Caption and NeSyGeo-CoT, comprising 100k samples.
-NeSyGeo-Caption aims to improve the perceptual understanding of geometric elements, while NeSyGeo-CoT primarily focuses on enhancing logical reasoning.
-The key characteristics of our dataset compared to other popular multimodal geometric datasets are presented in Figure 2.
-Additionally, we develop an evaluation set, NeSyGeo-Test, with 2668 Q&A pairs, enabling a thorough assessment of the geometric reasoning capabilities of mainstream MLLMs.
-We conducted an extensive and comprehensive evaluation of the geometric reasoning capabilities of current mainstream open-source and closed-source models, with details presented in Appendix 8.
-Notably, our training dataset consistently and efficiently enhances the geometric reasoning performance of MLLMs across multiple benchmarks.
-With only 4k samples and two epochs of RL training, base models achieve performance improvements of up to +15.8% on MathVision, +8.4% on MathVerse, and +7.3% on GeoQA.
-Moreover, InternVL2.5-4B can be improved to outperform the 8B model in the same series on geometric reasoning tasks.
+More specific examples of different methods are in Appendix E.1.
 In summary, our contributions are as follows:
-  • We propose NeSyGeo, a novel framework for geometric reasoning data generation, featuring a Geo-DSL for symbolic synthesis, a conversion engine for image and text generation, and an LLM-driven generator for Q&A pairs with CoT.
-NeSyGeo ensures validity through rigorous symbolic definitions and diversity via varied actions and neural searching.
-  • Using our framework, we synthesize the NeSyGeo-Caption and NeSyGeo-CoT training datasets with 100k high-quality samples, alongside a comprehensive geometric task evaluation set NeSyGeo-Test.
+  • We propose NeSyGeo, a novel neuro-symbolic framework for geometric reasoning data generation.
+NeSyGeo incorporates a comprehensive Geo-DSL, a reasoner–verifier for CoT generation, and a painter–translator module for informalization.
+It ensures data quality, enhances diversity, improve the correctness of reasoning chains, and reduces redundancy across modalities.
+  • Using our framework, we synthesize NeSyGeo-Caption and NeSyGeo-CoT training datasets with 100k samples, featuring high-quality images and reasoning chains, alongside a comprehensive geometric task evaluation set NeSyGeo-Test with 2,668 samples.
 These datasets are characterized by their diversity, rigor, and balanced distribution of information across image and text modalities.
@@
 Early approaches to geometric reasoning predominantly relied on symbolic solvers that used formal languages to tackle the tasks.
-For instance, Inter-GPS [18] and PGDP [35] employed symbolic methods by manually crafting reasoning rules and symbolic representations for geometric entities.
+For instance, Inter-GPS (Lu et al., 2021) and PGDP (Zhang et al., 2022) employed symbolic methods by manually crafting reasoning rules and symbolic representations for geometric entities.
 These systems typically transform visual input into symbolic forms through instance segmentation and apply theorem search to derive solutions.
 However, these methods lacked scalability due to their dependence on manually designed rules.
-Their inability to generalize beyond specific problem types further limited their universality and effectiveness across diverse geometric challenges.
-Figure 3: The overview of our neuro-symbolic data generation framework.
-The framework comprises three steps: In the first step, we get a symbolic language sequence in a limited symbolic action space.
-In the second step, the conversion engine parses the Geo-DSL sequence and translates it back to natural language and visual image without losing soundness.
-In the third step, we employ LLMs to take a reverse search and forward validation process to get final Q&A pairs with CoT.
+Their inability to generalize beyond specific problem types further limited their universality across diverse geometric challenges.
 The advent of MLLMs has shifted the paradigm toward data-driven geometric reasoning, leveraging their robust reasoning capabilities.
-Recent advancements include GeoDRL [19]and GeoGen [13].
-Despite these developments, geometric reasoning poses significant challenges for MLLMs, requiring seamless integration of image perception, geometric knowledge, and multi-step reasoning.
-GeoSense [29] identifies the identification and application of geometric principles as a persistent bottleneck.
-Similarly, GeoEval [34] reveals that current MLLMs exhibit significantly low accuracy when facing more challenging geometric problems.
-MathVerse [36] further highlights MLLMs’ over-reliance on textual information, underscoring the critical need for balanced multimodal datasets to enhance cross-modal reasoning capabilities.
+Recent advancements include GeoDRL (Peng et al., 2023) and GeoGen (Krueger et al., 2021).
+Despite these developments, geometric reasoning poses significant challenges for MLLMs, requiring image perception, geometric knowledge, and multi-step reasoning.
+GeoSense (Xu et al., 2025) identifies the identification and application of geometric principles as a persistent bottleneck.
+Similarly, GeoEval (Zhang et al., 2024b) reveals that current MLLMs exhibit significantly low accuracy when facing more challenging geometric problems.
+MathVerse (Zhang et al., 2024c) further highlights MLLMs’ over-reliance on textual information, underscoring the critical need for balanced multimodal datasets to enhance cross-modal reasoning capabilities.
 2.2 Multimodal Geometry Datasets
 Large-scale, high-quality datasets are essential for enhancing the performance of MLLMs in solving geometric problems.
-Early datasets such as GeoS [21] (186 problems), Geometry3k [18] (3000 problems) and GeoQA [4] (4998 problems) utilized human manual annotation.
+Early datasets such as GeoS (Seo et al., 2015) (186 problems), Geometry3k (Lu et al., 2021) (3000 problems) and GeoQA (Chen et al., 2021) (4998 problems) utilized human manual annotation.
 Their datasets are thus limited to a small scale.
@@
 To address this, numerous efforts have shifted toward automatic data generation.
-G-LLaVA [9] rephrased questions from GeoQA and Geometry3k to create 115,000 Q&A pairs, but failed to enhance image variety.
-Template-based methods [37, 7] typically rely on 10–20 predefined geometric figures, limiting the diversity of the generated images.
-AlphaGeometry [23], a notable work that combines symbolic solvers for geometric proofs, employs a symbolic language definition.
-Yet, due to the absence of numerical attributes such as angle measures and segment lengths in its geometric space, attempts to automatically generate datasets using the AlphaGeometry framework [10, 32] are confined to caption datasets, failing to produce the numerical Q&A pairs critical for current MLLMs training.
+Figure 3: Our pipeline is centered around a symbolic language Geo-DSL.
+The Generator synthesizes a sequence in this language, from which Reasoner and Verifier produce logically sound Q&A pairs and reasoning chains.
+Subsequently, Painter and Translator render the symbolic core into semantically orthogonal visual and textual outputs.
+G-LLaVA (Gao et al., 2025) rephrased questions from GeoQA and Geometry3k to create 115,000 Q&A pairs, but failed to enhance image variety.
+Template-based methods (Zhang et al., 2024d; Deng et al., 2024) typically rely on 10–20 predefined geometric figures, limiting the diversity of the generated images.
+AlphaGeometry (Trinh et al., 2024), a notable work that combines symbolic solvers for geometric proofs, employs a symbolic language definition.
+Yet, due to the absence of numerical attributes such as angle measures and segment lengths in its geometric space, attempts to automatically generate datasets using the AlphaGeometry framework.On the other hand,  (Huang et al., 2025) and  (Zhang et al., 2024a) are confined to caption datasets, failing to produce the numerical Q&A pairs critical for current MLLMs training.
 In contrast to prior approaches, our method pioneers a neuro-symbolic framework, being the first to integrate the precision of symbolic definition with the diversity of neural search for generating multimodal reasoning data.
 3 Methods
-To address the urgent need for large-scale and high-quality multimodal datasets in MLLMs for geometric reasoning, we propose NeSyGeo, a novel three-stage data generation pipeline.
-The pipeline is built upon Geo-DSL, a symbolic DSL designed to represent most elements in plane geometry space concisely and wholly.
-Its entity-relation-constraint structure allows any element to
-Table 1: Examples of our Geo-DSL and Corresponding Natural Language.
-Geo-DSL is defined by an entity-relation-constraint framework, encompassing 13 point types, 7 line types, 3 angle types, and 14 shape types in plane geometry.
-See Appendix H for complete Geo-DSL definitions.
-Type Geo-DSL Language Natural Language
-Shape $Triangle(A,B,C)=(x,y,\alpha)$ Triangle $ABC$ has $AB=x$, $BC=y$, $\angle B=\alpha$
-$Circle(O)=(x)$ Circle $O$ has radius $x$
-Point $Foot(D,A,Line(B,C))$ $D$ is the foot of the perpendicular from $A$ to $BC$
-$Intersection(E,Line(A,B),Line(C,D))$ $E$ is the intersection of line $AB$ and line $CD$
-Line $Para(Line(A,B),Line(C,D),x)$ Line $AB$ is parallel to $CD$, $AB=x$
-Angle $Angle(P,Q,R)=\alpha$ $\angle PQR=\alpha$
-be defined via a single statement, while its expressive power ensures comprehensive coverage of all geometric elements and values.
-NeSyGeo’s generation process unfolds in three distinct stages: First, a symbolic generator performs action augmentations within the finite symbolic space to produce a Geo-DSL sequence.
+The overall framework of NeSyGeo is illustrated in Figure 3 .The pipeline is built upon Geo-DSL(Section 3.1), with generation process unfolds in three parts:
+First, a symbolic generator performs action augmentations within the finite symbolic space to produce a Geo-DSL sequence.
 This approach effectively constrains the synthesis and augmentation process, ensuring validity and controllability by generating outside the infinite domains of natural language and image spaces (Section 3.2).
-Second, a conversion engine maps the generated Geo-DSL sequences back into natural language descriptions and visual image representations.
+Second, to get questions with reasoning paths, we utilize Reasoner and Verifier with a two-stage generation process.
+The reverse search primarily ensures the diversity of the Q&A pairs, while the forward verification confirms the correctness of the CoT and the final answer.
+(Section 3.3)
+Third, painter and translator map the generated Geo-DSL sequences back into natural language descriptions and visual image respectively.
 This process synthesizes high-quality images and valid text while avoiding intermodal information overlap.
-Third, to get Q&A Pairs with reasoning paths, we utilize expert LLMs to conduct backwards search to identify the geometric unknowns to be solved and generate the CoT in a forward manner (Section 3.4).
-The search process primarily ensures the diversity of the Q&A pairs, while the forward verification confirms the correctness of the CoT and the final answer.
-The overall framework of NeSyGeo is illustrated in Figure 3.
+(Section 3.4)
 3.1 Symbolic Definition
-Existing symbolic languages for plane geometry have certain limitations.
-The definitions in AlphaGeometry [23] are tailored for proof-based problems, thus lacking definitions related to specific numerical values.
-InterGPS [18] defines a predicate as a geometric shape entity, geometric relation, or arithmetic function, constructing 91 predicates.
-However, this approach overly fragments shape, attribute, and relation definitions into independent statements, often requiring multiple statements to specify a single element in the figure.
-This significantly increases the complexity of a conversion engine’s identification of elements within the symbolic space and further conversion of them.
-We propose Geo-DSL, a concise and comprehensive symbolic language for plane geometry to address these limitations.
-It employs an entity-relation-constraint framework, using well-defined rules to uniquely define 13 types of points, 7 types of lines, 3 types of angles, and 14 types of shapes, covering all plane geometry elements and incorporating numerical attributes like lengths and angles for precise specifications.
-Table 1 provides partial examples of symbolic definitions.
+Existing symbolic languages present significant limitations.
+The syntax of AlphaGeometry (Trinh et al., 2024), for example, is tailored for formal proofs and omits numerical specifications.
+In contrast, InterGPS (Lu et al., 2021) suffers from excessive fragmentation, requiring multiple statements to define single elements, which substantially complicates the parsing and conversion process.
+To surmount these challenges, we introduce Geo-DSL, a concise and comprehensive symbolic language for plane geometry.
+It is architected around an entity–attributes-relations framework, which we denote as a triplet $\mathcal{G}=(\mathcal{E},\mathcal{A},\mathcal{R})$, representing entities, attributes, and relations respectively.
+The syntax supports three primary declarative constructs: (i) Attribute Assertions ($e\rightarrow A_{e}$), which bind quantitative properties to an entity; (ii) Relational Assertions ($(e_{1},\dots,e_{n})\rightarrow r$), which define relationships among entities; and (iii) Hybrid Declarations that compose the two.
 Geo-DSL offers two key advantages.
-First, its comprehensive coverage includes all geometric elements and their numerical properties, enabling accurate and complete descriptions.
-Second, its simplicity allows a single statement to specify an element uniquely, promoting the incremental integration of the symbolic action space and the sequential parsing of statements by the conversion engine.
-This combination of completeness and simplicity makes Geo-DSL an efficient and powerful geometric representation and processing solution.
+First, it achieves descriptive completeness through a structured ontology that classifies all geometric primitives into four fundamental types (points, lines, angles, and shapes), representing the plane geometry with only 37 core statements.
+Second, its syntactic parsimony, where each entity is unambiguously instantiated via a single statement, enables both the principled composition of the symbolic action space and its deterministic parsing by our conversion engine.
+This synthesis of expressive power and structural minimalism renders Geo-DSL a robust and scalable framework for automated geometric representation and processing.
 3.2 Symbolic Sequence Generation
-To generate a Geo-DSL sequence, we introduce a step-action augmenter that iteratively synthesizes a sequence of statements, as detailed in Algorithm 1.
-Based on dataset preferences, we first configure the step count $N$, weight matrices $I$ and $A$ for selecting elements and actions with respective probabilities, and ranges $[l_{\text{min}},l_{\text{max}}]$ for lengths and $[\theta_{\text{min}},\theta_{\text{max}}]$ for angles.
-Then, the augmenter iteratively generates symbolic statements over $N$ steps.
-For each step, we randomly sample parameters $x,y,z$ and $\alpha$, select an element $v_{j}\in f_{v}$ using weights from $I$, and choose an action $a_{k}$ based on weights from $A$ (see Appendix I for action details).
-The new statement $s_{\text{new}}$ is then incorporated into the sequence $f_{s}$.
-Leveraging Geo-DSL’s symbolic definitions and well-defined actions, this approach ensures the validity and accuracy of each statement.
-Meanwhile, randomized elements, diverse action selections, and customizable hyperparameter preferences promote diversity.
-Algorithm 1 The overall framework of the symbolic sequence generation process
-1:Input: Step count $N$, Action weight matrix $A$, Element selection weight matrix $I$, Line length range $[l_{min},l_{max}]$, Angle range $[\theta_{min},\theta_{max}]$$\triangleright$ Set customizable hyperparameter
-2:Output: Generated Geo-DSL statement sequence $f_{s}$.
-3:Initialize $f_{s}$=Initialize( ) . $\triangleright$ Initialize the sequence $f_{s}$ with the first statement
-4:Initialize symbolic state space elements $f_{v}=\text{Initialize}(f_{s})$. $\triangleright$ Initialize state based on $f_{s}$
-5:for $i=1$ to $N$ do
-6:     Randomly sample $x,y,z$ from $[l_{min},l_{max}]$.
-7:     Randomly sample $\alpha$ from $[\theta_{min},\theta_{max}]$.
-8:     Element $v_{j}$=Selected_elements($f_{v}$,$I$). $\triangleright$ Select element $v_{j}$ from $f_{v}$ randomly.
-9:     Action $a_{k}$=Selected_action($v_{j}$,$A$). $\triangleright$ Select action $a_{k}$ based on the type of $v_{j}$ randomly
-10:     $s_{new}$=Generate_DSL($a_{k}$, x, y, z, $\alpha$). $\triangleright$ Generate new DSL statement
-11:     $f_{s}$=Update($f_{s}$,$s_{new}$) $\triangleright$ Add the new statement to the sequence
-12:     $f_{v}$=Update($f_{v}$,$s_{new}$). $\triangleright$ Update state space elements
-13:end for
-14:return $f_{s}$.
+Generator: To generate a Geo-DSL sequence, we employ a step-action augmentation procedure that iteratively synthesizes a sequence of statements, as detailed in Algorithm 1.
+Our methodology is initialized by configuring hyperparameters based on dataset characteristics: the sequence length $N$, weight matrices $\mathbf{I}$ and $\mathbf{A}$ that parameterize the selection distributions for geometric elements and symbolic actions (see Appendix H.2 for action details), and the value ranges $l_{\text{min}},l_{\text{max}}]$ and $\theta_{\text{min}},\theta_{\text{max}}]$ for lengths and angles, respectively.
+The augmenter then commences an iterative synthesis over $N$ steps.
+At each step, after sampling continuous parameters ($x,y,z,\alpha$) and assigning randomized labels to the geometric vertices, our generative process selects a discrete element-action pair $(v_{j},a_{k})$ using a hierarchical weighted sampling strategy.
+First, an element $v_{j}$ is sampled from the available set $f_{v}$.
+The probability is proportional to its weight from the matrix $\mathbf{I}$, modulated by an element selection temperature $\tau_{e}$:
+$P(v_{j})=\frac{(\mathbf{w}_{I}(v_{j}))^{1/\tau_{e}}}{\sum_{j^{\prime}\in f_{v}}(\mathbf{w}_{I}(v_{j^{\prime}}))^{1/\tau_{e}}}$ (1)
+Subsequently, conditioned on the chosen element $v_{j}$, a compatible action $a_{k}$ is sampled from the valid set $\mathcal{A}(v_{j})$, with its probability determined by weights from the matrix $\mathbf{A}$ and modulated by an action selection temperature $\tau_{a}$:
+$P(a_{k}|v_{j})=\frac{(\mathbf{w}_{A}(a_{k},v_{j}))^{1/\tau_{a}}}{\sum_{k^{\prime}\in\mathcal{A}(v_{j})}(\mathbf{w}_{A}(a_{k^{\prime}},v_{j}))^{1/\tau_{a}}}$ (2)
+The newly formed statement $s_{\text{new}}$ is then integrated into the sequence $f_{s}$.
+This two-stage, temperature-controlled probabilistic approach guarantees the validity of each statement while promoting significant diversity in the generated sequences.
+3.3 CoT Generation
+To generate Q&A pairs accompanied by reasoning paths, we first formalize the underlying deductive task.
+The objective is to navigate from the initial premises to a valid conclusion through a sequence of logical steps.
+This process can be conceptualized through the formulation of a reasoning graph:
+$\mathcal{G}=(\mathcal{S},\mathcal{R},\hookrightarrow)$
+where:
+  • $\mathcal{S}$ is the set of all possible states, where each state $s\in\mathcal{S}$ represents the initial state or a deduced conclusion within the reasoning process.
+  • $\mathcal{R}$ denotes the set of geometric rules, with each rule $r\in\mathcal{R}$ encoding a principle of logical deduction that governs the relationships between states.
+  • $\hookrightarrow\subseteq\mathcal{S}\times\mathcal{R}\times\mathcal{S}$ is the state transition relation.
+A transition $(S_{r},r,s^{\prime})$ signifies that state $s^{\prime}$ is derived from a subset of states $S_{r}\subset\mathcal{S}$ by applying rule $r$.
+However, a direct implementation of this formalism via a deterministic inference engine is fraught with significant limitations.
+Firstly, such an engine is computationally prohibitive, as it necessitates an exhaustive enumeration of the combinatorial space of conditions at each intermediate state.
+Given a state characterized by a set of $N$ known conditions, the engine must evaluate the applicability of $C$ fundamental axioms and theorems.
+For each rule $i$, which is predicated on matching a specific set of $c_{i}$ conditions, the search space grows combinatorially, with a complexity lower-bounded by $\mathcal{O}\left(\sum_{i=1}^{C}\binom{N}{c_{i}}\right)$, rendering any exhaustive search intractable for non-trivial geometric configurations.
+Secondly, this approach exhibits profound domain specificity.
+The reasoning framework and its constituent logical rules are intrinsically hard-coded for plane geometry.
+This tight coupling to a single domain severely curtails the model’s generalizability and fundamentally impedes its application for synthesizing datasets in other domains without substantial, non-trivial re-engineering.
+To overcome these limitations, we pivot to leveraging the capabilities of LLMs, which can harness their extensive pre-trained knowledge and implicit intuitive faculties to facilitate rapid node expansion within the reasoning graph, obviating the need for bespoke rule engineering and thus offering a more agile and convenient paradigm.
+However, modern LLMs still face challenges when tackling intricate geometric tasks, and employing them for a direct, one-shot generation of Q&A pairs can result in erroneous information.
+To mitigate these issues, we introduce a novel two-stage methodology, which comprises a progressive reverse-search reasoner and a forward-validation verifier.
+Reasoner.
+We introduce a novel search paradigm that operationalizes deductive reasoning as a Breadth-First Search (BFS) over the deductive lattice, accelerated by the learned preferences of LLMs.
+Instead of a uniform exploration, we leverage the reasoner to steer the expansion of the search frontier.
+From any given state $S_{t}$, the model $M_{\theta}$ is prompted to infer a single, contingent, and computable geometric conclusion $c_{t}$.
+This new conclusion augments the current state, effecting a state transition defined by $S_{t+1}=S_{t}\cup\{c_{t}\}$, where $c_{t}\leftarrow M_{\theta}(S_{t})$.
+This iterative expansion establishes a form of progressive supervision and effectively prunes the vast search space, ensuring the fidelity and diversity of the resultant reasoning trajectories.
+By decomposing the complex, long-horizon inference task into a sequence of incremental, verifiable deductions, it substantially lowers the reasoning complexity and mitigates the model’s propensity for hallucination.
+The terminal conclusions derived from this process are then randomly sampled to form candidate Q&A pairs.
+Verifier.
+To verify the logical integrity of the generated Q&A pairs and to articulate a step-by-step CoT reasoning process.
+The questions generated by Reasoner, along with the corresponding conditions, are supplied as input.
+Verifier is prompted to produce both a final answer and its reasoning trajectory.
+These outputs are then cross-referenced against the answers derived from Reasoner.
+Only those pairs exhibiting concordance are retained for inclusion in our final dataset.
+The dual-verification protocol not only guarantees the correctness of the answers but also yields diverse and valid CoT narratives without necessitating an exhaustive traversal of the entire reasoning space.
 3.4 Informalization
-Following the generation of sequences within the formal symbolic space, it is necessary to map them back to the text and image spaces for further processing and visualization.
-Our approach generates high-quality images and rigorous natural language, allocating information between them to compel MLLMs to leverage visual data effectively.
-Figure 4: Reverse search and forward validation with expert LLMs
-Visual image.
-For the visual space, our visualization engine parses each Geo-DSL statement as a geometric element to generate high-quality images using Matplotlib.
-The rigorous symbolic language space ensures the determinacy and uniqueness of the conversion process, thereby enabling precise image generation.
-Additionally, the engine produces images with detailed annotations absent from the text, requiring models to leverage images during problem-solving, enhancing their visual perception and ability to extract image-based information.
-Natural language.
-We adopt a template-based transformation approach, parsing and mapping each symbolic statement to multiple predefined natural language templates.
-This produces a text-full version containing all details and captions of the image and a text-lite version retaining only essential conditions, not included in the image annotations, as the final form in our datasets.
-The text-lite version avoids intermodal information redundancy, while diverse templates ensure the validity and diversity of the synthesized condition text.
-3.4 CoT Generation
-To generate Q&A pairs with CoT reasoning, we utilize the strong reasoning abilities of expert LLMs to ensure diverse and reliable output.
-Using the text-full version generated as described in Section 3.3 as input, we develop a two-step process comprising reverse search and forward validation as shown in Figure 4.
-Prompts are detailed in Appendix G.
-Reverse search.
-We use DeepSeek R1 [6] for reverse search, starting from the given conditions, iteratively exploring and deriving conclusions step by step, and ultimately producing the final conclusions at the end of the reasoning chain along with their corresponding answers.
-This reverse approach starts from known information and recursively builds a reasoning chain, reducing Q&A generation complexity and hallucinations.
-R1’s strong reasoning and exploration capabilities yield diverse conclusions, enriching the variety of Q&A pairs.
-Forward validation.
-To ensure the correctness of Q&A pairs and generate step-by-step CoT reasoning, we re-input the questions and text-full conditions into DeepSeek V3, requesting both the reasoning process with final answer, and cross-validate these answers against those from R1 to include only consistent pairs in our final dataset.
-This process guarantees answer correctness and yields diverse, valid CoT without an extensive search space.
+The formal symbolic sequences from our Geo-DSL must be mapped back into the visual and textual modalities.
+To compel cross-modal reasoning, we enforce a principle of information orthogonality: All explicit constraints and premises $I$ in initial state $S_{0}$are partitioned into disjoint visual modality $I_{v}$ and textual modality $I_{t}$ , such that $I_{v}\cap I_{t}=\emptyset$.
+Painter.
+For the visual modality, painter parses each Geo-DSL statement as a geometric primitive, deterministically rendering high-fidelity images via Matplotlib.
+The rigorous symbolic space ensures the determinacy and uniqueness of the conversion process.
+Crucially, to mitigate inter-modal information imbalance, Painter employs a visual-centric annotation strategy, which involves embedding detailed information $I_{v}$ within the image space that are explicitly withheld from the accompanying textual description.
+Please refer to  A.2 for more details.
+This design choice robustly compels MLLMs to intrinsically leverage visual cues during problem-solving, thereby fostering enhanced visual perception and facilitating the extraction of perceptually-grounded information.
+Translator.
+Translator converts each symbolic statement to multiple predefined natural language templates.
+To ensure diversity and randomness, each template is stochastically generated by GPT-4o and subsequently undergoes a manual curation process.
+This guarantees that every geometric element is associated with a minimum of eight corresponding descriptive templates.
+The final text condition avoids intermodal information redundancy, while diverse templates ensure the validity and diversity of the synthesized condition text.
 4 Experiments
-We present a series of experiments to investigate the following four research questions:
-Efficacy – To what extent does training on our synthesized dataset in both RL and SFT improve the geometric reasoning performances of several MLLMs?
-Efficiency – Is the data generated by NeSyGeo more effective in yielding models with better performance compared to using data from existing automatic synthesis frameworks?
-Diversity – Does the NeSyGeo framework effectively ensure diversity across both the text and image spaces of the generated synthetic geometric dataset?
-Visual Effectiveness – Can our datasets compel models to effectively utilize visual information for enhanced understanding by appropriately distributing information between modalities?
 4.1 Experimental Setup
 Dataset.
-To synthesize a diverse dataset, we set hyperparameters for generating images by configuring the step count, length range, and angle range.
-The step count $N$ ranges from one to four.
-The line length is defined within the basic range $[l_{\text{min}},l_{\text{max}}]=[1,5]$, scalable by any multiple of 2 (e.g., $[4,20]$).
-The angle is constrained to multiples of 15° within $[15^{\circ},165^{\circ}]$, with increased weights assigned to special angles.
-We generate various types of weight matrices $A$ and $I$ by adjusting their corresponding values.
-This process yields a NeSyGeo-CoT dataset with 30k Q&A pairs and a NeSyGeo-Caption dataset with 70k Q&A pairs.
-Additional dataset statistics are provided in Appendix B.
+We generate a NeSyGeo-CoT dataset with 30k Q&A pairs and a NeSyGeo-Caption dataset with 70k Q&A pairs.
+The settings of generation process can be found in Appendix C.1.
+Additional dataset statistics and features are provided in Appendix B.1.
 Evaluation.
-Our evaluation is conducted on several benchmarks: the Test set of GeoQA[4], the Test_MINI set of MathVision[26], and the MathVerse[36].
-For the MathVerse benchmark, we select the Vision Only, Vision Dominant, and Vision Intensive sets to better assess the visual perception and logical reasoning capabilities of MLLMs.
+Our evaluation is conducted on several benchmarks: the Test set of GeoQA(Chen et al., 2021), the Test MINI set of MathVision(Wang et al., 2024b), and the MathVerse(Zhang et al., 2024c).
+For the MathVerse benchmark, we select the Vision Only, Vision Dominant, and Vision Intensive sets to better assess the visual perception and logical reasoning capabilities.
 We extract in-domain metrics from other datasets, including angle, area, length, and Plane Geometry, to effectively evaluate the models’ capabilities in geometric reasoning problems, in addition to the GeoQA dataset, which focuses entirely on plane geometry.
-For GeoQA, we employed hard-coded extraction for comparison, while other evaluations are assessed using the automated VLMEvalKit framework[8].
+For GeoQA, we employed hard-coded extraction for comparison, while other evaluations are assessed using the automated VLMEvalKit framework(Duan et al., 2024).
 Appendix C provides additional experimental details and evaluation results.
 4.2 Empirical Results
-Efficacy: Training multiple MLLMs with NeSyGeo via both SFT and RL significantly enhances geometric problem-solving performances.
+Quality: Manually evaluated by human experts, our dataset was found to contain high-quality CoT rationales, accurate Q&A pairs, and aesthetically pleasing images.
+Figure 4: Human evaluation results comparison.
+To assess our dataset, we conducted a human expert evaluation with 40 experts (all at university undergraduate level or higher), comparing our NeSyGeo-CoT against G-LLaVA (Gao et al., 2025) and MAVIS(Zhang et al., 2024d).
+The evaluation was structured around five dimensions comprising nine metrics: Image, Question (including conditions), Reasoning path, Overall Impression and Final Answer.
+Each metric was rated on a 5-point scale.
+More details can be found in appendix C.2.
+NeSyGeo achieves state-of-the-art results on most metrics, ensuring a high degree of correctness, quality, and aesthetics, as corroborated by human evaluation.
+To validate our two-stage generation strategy on CoT and Q&A quality, we also performed an ablation study on the Reasoner and Verifier.
+As shown in Table 1, we generated 50 instances for each configuration.
+The results demonstrate that the Reasoner successfully expands the deductive search space and reduces reasoning hallucinations, which ultimately enhances both the problem difficulty and the depth of the reasoning chains.
+Concurrently, the Verifier provides effective supervision that is instrumental in ensuring the logical soundness and high accuracy of the final data.
+Table 1: Ablation study on the effects of the Reasoner and Verifier.
+Pass Rate measures the proportion of solutions accepted by the Verifier, while Pass@1 Rate (gpt-4o-mini’s) serves as a proxy for problem difficulty.
+All other metrics align with our human preference experiments.
+Configurations Pass Rate CoT-Quality CoT-Accuracy Ans-Correctness Pass@1 Rate
+Reasoner Verifier
+X 78.0 4.34 4.66 5.00 68.0
+X 4.40 4.22 4.42 52.0
+64.0 4.68 4.54 5.00 48.0
+Performance: Our dataset enhances the geometric reasoning capabilities of models through both RL or SFT.
+Under the same data budget, ours is better than that from popular automatic generation frameworks.
 We first sample 4k samples from our NeSyGeo-CoT dataset and apply the Group Relative Policy Optimization (GRPO) algorithm to train two epochs with Deepseek R1’s format and answer rewards.
-The training code framework is based on VLM-R1 [22].
-As shown in Tables 2, models achieve the best performance among the baselines after training.
-InternVL2.5-4B significantly improved in the angle knowledge domain with gains of 8.4 (MathVerse), 7.3 (GeoQA), and 5.3 (MathVision).
-Qwen2.5-VL-3B achieved a +15.8 performance boost in the area domain of MathVision.
+As shown in Tables 2, InternVL2.5-4B significantly improved in the angle knowledge domain with gains of 8.4 (MathVerse), 7.3 (GeoQA), and 5.3 (MathVision).
+Qwen2.5-VL-3B achieved +15.8 performance boost in the area domain of MathVision.
 Notably, across all evaluated metrics, the InternVL2.5-4B model trained on the NeSyGeo dataset achieves performance on par with or superior to its 8B counterpart.
-We also conducted SFT experiments, initially training on our NeSyGeo-Caption dataset to enhance the models’ perception of geometric images, followed by training on the NeSyGeo-CoT dataset to improve reasoning capabilities.
-The experiments were conducted on LLaMA-Factory [40] framework.
-Evaluation results on MathVerse (Vision Intensive) and GeoQA are presented in Table 3.
-The trained model demonstrates performance improvements over the base model on most metrics.
-Efficiency: Under the same data budget, the generated data from our framework is better than that from popular automatic generation frameworks.
-We randomly sampled 4k samples from MAVIS [37] and R-CoT [7], which are automatic frameworks in geometry problem generation. To ensure a fair comparison, we maintained consistent settings.
+To ensure a fair comparison with others, we randomly sampled same budget from MAVIS (Zhang et al., 2024d) and R-CoT (Deng et al., 2024) and maintained consistent settings.
+As illustrated in Figure 5 and Table 2, while all datasets help the model improve over the baseline, our dataset outperformed others in most metrics, validating its superior performance.
 Table 2: RL performance comparison: Models trained with only 4k samples of NeSyGeo-CoT show performance gains over the base models, with the InternVL2.5-4B model exceeding the 8B variant in geometry problem-solving.
 GeoQA MathVision MathVerse
-Model Angle Area Length Angle Area Length Plane Geometry
+Model Angle Area Length Angle Area Length Plane
 Qwen2.5-VL-3B 53.3 26.3 26.3 21.1 31.3 20.9 37.0 32.5
-Qwen2.5-VL-3B+NeSyGeo 55.7 (+2.4) 26.3 (+0.0) 42.1 (+15.8) 26.3 (+5.2) 32.6 (+1.3) 23.5 (+2.6) 37.2 (+0.2) 35.5 (+3.0)
+Qwen2.5-VL-3B+Ours 55.7 (+2.4) 26.3 (+0.0) 42.1 (+15.8) 26.3 (+5.2) 32.6 (+1.3) 23.5 (+2.6) 37.2 (+0.2) 35.5 (+3.0)
 InternVL2.5-4B 61.9 36.8 31.6 26.3 31.5 22.7 31.9 30.7
@@
 InternVL2.5-4B+R-CoT 63.3 (+1.4) 31.6 (-5.2) 31.6 (+0.0) 21.1 (-5.2) 31.2 (-0.3) 18.3 (-4.4) 34.3 (+2.4) 28.7 (-2.0)
-InternVL2.5-4B+NeSyGeo 69.2 (+7.3) 42.1 (+5.3) 36.8 (+5.2) 26.3 (+0.0) 39.9 (+8.4) 24.9 (+2.2) 36.1 (+4.2) 36.7 (+6.0)
+InternVL2.5-4B+Ours 69.2 (+7.3) 42.1 (+5.3) 36.8 (+5.2) 26.3 (+0.0) 39.9 (+8.4) 24.9 (+2.2) 36.1 (+4.2) 36.7 (+6.0)
 InternVL2.5-8B 66.2 36.8 36.8 21.1 36.9 23.1 34.8 36.6
 Table 3: SFT performance comparison: The trained model demonstrates performance improvements over the base model on most metrics.
-GeoQA Vision Intensive
-Model Angle Area Length Plane Geometry
+GeoQA MathVerse
+Model Angle Area Length Plane
 Qwen2.5-VL-7B 69.4 43.0 27.5 46.2 44.1
-Qwen2.5-VL-7B+NeSyGeo 71.8 (+2.4) 46.1 (+3.1) 23.1 (-4.4) 49.5 (+3.3) 46.7 (+2.6)
-LLaVA-Next-7B 22.6 28.5 6.6 16.5 20.4
-LLaVA-Next-7B+NeSyGeo 26.1 (+3.5) 30.6 (+2.1) 7.7 (+1.1) 19.2 (+2.7) 22.9 (+2.5)
-As illustrated in Figure 5 and Table 2, while all datasets help the model improve over the baseline, our dataset outperformed others in most metrics, validating its superior performance.
-Diversity: The dataset synthesized by our NeSyGeo framework exhibits high diversity in both text and visual features.
-A critical challenge for automatic data synthesis methods is whether the dataset is sufficiently diverse to avoid quality degradation due to potential overfitting risks from inherent domain constraints.
-We employ t-SNE [24] dimensionality reduction for mapping in text space to evaluate the diversity across different methods.
-This analysis allows us to assess the diversity of the textual descriptions themselves.
-Given that in geometric problems, the text conditions depict specific visual elements, and the diversity observed in the text space also serves as a valuable indicator of the diversity in the corresponding visual diagrams.
-To ensure a fair comparison, we remove all prompts related to guiding large models, retaining only condition and question texts, and randomly sample 5k texts from each dataset.
-The results are illustrated in Figure 6.
-Our method and G-LLaVA [9] exhibit uniformly distributed features in the space, indicating low data overlap and high diversity.
-In contrast, R-CoT and MAVIS display varying degrees of clustered distribution, indicating more feature-similar samples.
-To directly assess the diversity of the visual features, we also performed t-SNE on image features extracted by ResNet, with detailed experimental results presented in Appendix C.
+Qwen2.5-VL-7B+Ours 71.8 (+2.4) 46.1 (+3.1) 23.1 (-4.4) 49.5 (+3.3) 46.7 (+2.6)
+LLaVA-7B 22.6 28.5 6.6 16.5 20.4
+LLaVA-7B+Ours 26.1 (+3.5) 30.6 (+2.1) 7.7 (+1.1) 19.2 (+2.7) 22.9 (+2.5)
+We also conducted SFT experiments, initially training on our NeSyGeo-Caption dataset to enhance the models’ perception of geometric images, followed by training on the NeSyGeo-CoT dataset to improve reasoning capabilities.
+Evaluation results on MathVerse (Vision Intensive) and GeoQA are presented in Table 3.
+The trained model demonstrates performance improvements over the base model on most metrics.
 Figure 5: Efficiency comparison of our NeSyGeo-CoT dataset versus other mainstream automated synthesis datasets. The models are trained using RL methods with InternVL2.5-4B.
-Figure 6: T-SNE of the text features of different automatic frameworks.
-The G-LLaVA method augments the text space on the manually annotated GeoQA dataset.
-Thus, its text diversity can approximate that of real data more closely.
-Similar to G-LLaVA, our method exhibits a uniform distribution in the space, demonstrating superior diversity.
+Generability: The performance gains of models are from true reasoning, not just distilling solutions leaked from contaminated data
+Although we have implemented stringent safeguards to ensure that our datasets do not access the test set, a critical inquiry is whether the efficacy of our framework stems from inadvertent data leakage introduced during the generation process.
+Adopting the methodology delineated by Wu et al.
+(Wu et al., 2025), we conducted an experiment utilizing the GeoQA test set.
+For each question, we furnished the models with partial prompts, specifically truncated to the initial 40%, 60%, and 80% of their original length.
+We then evaluated two key metrics: (1) The ROUGE-L score, quantifying the textual overlap between the model’s generated output and gpt-4o’s output, and (2) The final answer accuracy.
+This empirical evaluation was performed on two distinct models, Qwen2.5-VL-3B and InternVL2.5-4B.
+While a marginal increase in certain ROUGE-L scores as shown in Table 4, we posit that this is a manifestation of the model’s augmented reasoning capabilities, which empower it to perform more structured and relevant explorations from a limited context.
+The consistently negligible accuracy scores across all truncated inputs strongly indicate that the models have not merely memorized the test instance.
+Consequently, we conclude that there is no substantive evidence to suggest that the model’s performance gains are attributable to data contamination.
+Table 4: Analysis of Potential Data Contamination using Truncated Prompts.
+0.4 Ratio 0.6 Ratio 0.8 Ratio
+Model ROUGE-L Acc. (%) ROUGE-L Acc. (%) ROUGE-L Acc. (%)
+QwenVL-base 28.04 0.00 29.88 0.00 31.95 0.00
+QwenVL-RL 21.29 0.00 36.72 0.00 40.57 0.00
+InternVL-Base 26.82 0.00 40.04 0.00 48.07 4.69
+InternVL-RL 30.00 0.00 44.91 0.00 57.37 1.54
+Visual Perception: Models trained on NeSyGeo data shows modest gains over information-redundant datasets, yet achieves substantial improvements when image understanding is necessary.
+This indicates that ours enhance model’s cross-modal reasoning.
+A key question is whether reducing redundancy and forcing models to extract visual information improves their geometric reasoning capabilities.
+We evaluate models on the MathVerse(Text Dominant), which provides redundant text descriptions and implicit properties enabling reasoning without images, and the MathVerse(Vision Only) version, where all information is embedded entirely within the images.
+For this comparison, we selected two text-redundant datasets: NeSyGeo+RED, the original NeSyGeo-CoT dataset supplemented with textual equivalents of its image annotations, and the R-CoT dataset.
+Results presented in Table 5 show that our model outperforms the baseline on Text-Dominant but lags behind other datasets in some metrics.
+On Vision-Only, our model surpasses them across all metrics, demonstrating enhanced geometric reasoning and visual perception.
 Table 5: Comparison between NeSyGeo-CoT and text-redundant datasets with equivalent data budgets.
@@
 R-CoT 51.8 24.2 45.0 46.5 25.9 23.1 31.3 29.2
-Visual Perception: Models trained on NeSyGeo data shows modest gains over information-redundant datasets when textual shortcuts exist, yet achieves substantial improvements when image understanding is necessary.
-This indicates that ours enhances not only logical reasoning but also image perception and utilization of the model.
-A key question is whether reducing redundancy and forcing models to extract visual information improves their geometric reasoning capabilities.
-We evaluate models on the MathVerse(Text Dominant), which provides redundant text descriptions and implicit properties enabling reasoning without images, and the MathVerse(Vision Only) version, where all information is embedded entirely within the images.
-For this comparison, we selected two text-redundant datasets: NeSyGeo+RED, the original NeSyGeo-CoT dataset supplemented with textual equivalents of its image annotations, and the R-CoT dataset.
-Results presented in Table 4 show that our model outperforms the baseline on Text-Dominant but lags behind other datasets in some metrics.
-On Vision-Only, our model surpasses them across all metrics, demonstrating enhanced geometric reasoning and visual perception.
 5 Conclusion
-This paper introduces NeSyGeo, a neurosymbolic framework for automatically synthesizing multimodal geometric datasets.
-Our approach transforms the generation process into a controllable symbolic space using Geo-DSL, maps the symbolic representation back to image and natural language spaces via a conversion engine, and then utilizes LLMs for backwards search and forward solving to produce Q&A pairs.
-Using this framework, we construct the NeSyGeo-CoT and NeSyGeo-Caption datasets, totalling 100k samples.
+This paper introduces NeSyGeo, a neuro-symbolic framework for automatically synthesizing multimodal geometric datasets.
+Our approach transforms the generation process into a controllable symbolic space using Geo-DSL, utilizes Reasoner and Verifier for reverse search and forward validation to produce Q&A pairs, and then maps the symbolic representation back to image and natural language spaces via Painter and Translator.
+Using this framework, we construct NeSyGeo-CoT and NeSyGeo-Caption datasets, totalling 100k samples.
 We also propose NeSyGeo-Test, a comprehensive benchmark for evaluating MLLMs’ geometric reasoning capabilities.
 Our datasets significantly and consistently improve the reasoning abilities of multiple MLLMs through both SFT and RL.
-Future Work: We intend to extend NeSyGeo to other multimodal domains, such as analytical geometry and visual question answering.
-This extensibility will be achieved by defining new domain-specific languages, corresponding synthesis rules within the symbolic space, and tailored conversion engines.
-Furthermore, we plan to develop an automated symbolic solver capable of conducting search and validation directly within the symbolic space.
-This would remove reliance on LLMs, potentially reducing generation costs and ensuring complete correctness of the datasets.
-Appendix A Comparison with Specific Examples of Popular Geometry Datasets
-To facilitate comparison of dataset characteristics synthesized by our method and other popular approaches, we showcase a randomly selected example from NeSyGeo-CoT alongside each of the different approaches in Figure 7.
-Geometry-3K is a manually synthesized dataset, while the remaining approaches employ automatic generation techniques.
-To ensure a fair comparison, we standardize the text format by removing model-guiding prompts and appending options when present.
-Furthermore, we annotate each sample with image pixels and CoT word counts.
-Compared to other datasets, our dataset features clear, human-aesthetically pleasing images, high-quality step-by-step reasoning chains, symbolic form meta-information enabling subsequent image augmentation and mutation, and well-distributed conditional information between images and text.
-Additional examples of our NeSyGeo-CoT dataset can be found in the Appendix 11.
-Figure 7: Comparison NeSyGeo-CoT dataset with other Popular Geometry Datasets.
-Geometry-3K is a manually synthesized dataset, while the remaining approaches employ automatic generation techniques.
-Our dataset features clear, human-aesthetically pleasing images, high-quality step-by-step reasoning chains, symbolic form meta-information enabling subsequent image augmentation and mutation, and well-distributed conditional information between images and text.
-Appendix B Statistics of NeSyGeo-Caption and NeSyGeo-CoT
-Detailed numerical statistics and element distribution for the NeSyGeo-Caption and NeSyGeo-CoT datasets are presented in Table 6 and  6.
+Ethics Statement
+NeSyGeo datasets presented in this paper are generated entirely through a synthetic, neuro-symbolic framework.
+The core generation process does not involve any user data, personally identifiable information, or web-scraping.
+Human involvement was limited exclusively to a voluntary study for evaluating the final dataset’s quality.
+All data collected from these human experts was fully anonymized to ensure strict confidentiality and the detailed evaluation protocol is provided in Appendix C.2.
+This methodology mitigates concerns related to both data privacy and social bias.
+Given the highly specialized nature of this mathematical task, we assess the potential for misuse or negative societal impact to be minimal.
+We release our contributions to the research community to responsibly foster progress in automated scientific reasoning.
+Reproducibility Statement
+To ensure the reproducibility of our research, we provide a comprehensive description of our data generation methodology throughout the main paper and in the Appendix A.
+Each module of our synthesis pipeline is delineated with its corresponding parameters and implementation details.
+We commit to making the complete synthetic dataset publicly available to the research community upon publication to facilitate further investigation and development.
+Furthermore, all hyperparameters and key configurations for our model training and experiments are fully disclosed in the Appendix C.
+Appendix Contents
+Appendix A More Details of Methods
+A.1 Generator
+As outlined in Algorithm 1, for each step, we first generate three lengths, $x$, $y$, and $z$, along with an angle $\alpha$, sampled from predefined ranges.
+We then randomly selecting vertex labels.
+Subsequently, based on the weight matrices $A$ and $I$, we determine the specific statement to be selected.
+The chosen statement, paired with its corresponding numerical values, is then appended to the Geo-DSL sequence.
+Algorithm 1 The overall framework of the symbolic sequence generation process
+1:Input: Step count NN, Action weight matrix $A$, Element selection weight matrix $I$, Line length range $l_{min},l_{max}]$, Angle range $\theta_{min},\theta_{max}]$$\triangleright$ Set customizable hyperparameter
+2:Output: Generated Geo-DSL statement sequence $f_{s}$.
+3:Initialize $f_{s}$=Initialize( ) . $\triangleright$ Initialize the sequence $f_{s}$ with the first statement
+4:Initialize symbolic state space elements $f_{v}=\text{Initialize}(f_{s})$. $\triangleright$ Initialize state based on $f_{s}$
+5:for $i=1$ to NN do
+6:  Randomly sample $x,y,z$ from $l_{min},l_{max}]$.
+7:  Randomly sample $\alpha$ from $\theta_{min},\theta_{max}]$.
+8:  Element $v_{j}$=Selected_elements($f_{v}$,II). $\triangleright$ Select element $v_{j}$ from $f_{v}$ randomly.
+9:  Action $a_{k}$=Selected_action($v_{j}$,AA). $\triangleright$ Select action $a_{k}$ based on the type of $v_{j}$ randomly
+10:  $s_{new}$=Generate_DSL($a_{k}$, x, y, z, $\alpha$). $\triangleright$ Generate new DSL statement
+11:  $f_{s}$=Update($f_{s}$$s_{new}$) $\triangleright$ Add the new statement to the sequence
+12:  $f_{v}$=Update($f_{v}$$s_{new}$). $\triangleright$ Update state space elements
+13:end for
+14:return $f_{s}$.
+A.2 Painter
+A primary challenge in machine-based geometric analysis is the frequent ambiguity between visual diagrams and the text that describes them.
+This disconnect between the visual and textual data streams creates a significant impediment for models attempting to perform integrated, cross-modal reasoning.
+To address this, we developed Painter, a rendering protocol that systematically embeds explicit semantic information directly into the geometric figures themselves.
+Painter ensures that each visual element has a clear and unambiguous interpretation by applying a strict set of annotation rules. The strategies are as follows:
+  • Segment Length: Each line segment is explicitly labeled with its identifier and corresponding numerical length, with the value rounded to one decimal place.
+  • Angle Annotations: The measure of any non-right angle is directly inscribed within the angle’s arc. For right angles, the conventional square symbol is used exclusively.
+  • Circle Radius: The radius of a circle is represented by a dashed line from the center to a point on the circumference, clearly marked with the label “r=[length]’‘.
+To further enhance the visual diversity and aesthetic quality of the generated figures, Painter incorporates several stochastic rendering strategies.
+For each synthesis, the global scale ( the pixel-to-unit ratio) is randomly sampled, and the geometric figure is randomly translated within the canvas.
+Additionally, distinct classes of geometric elements are rendered using a varied color palette to improve their visual discriminability.
+These procedures ensure a high degree of variation in the training data, preventing the model from overfitting to superficial stylistic features such as position or scale.
+Crucially, to prevent information redundancy and compel a true synthesis of visual and textual information, any property that is graphically rendered by Painter is intentionally omitted from the accompanying text description.
+A.3 Translator
+To provide the Reasoner with sufficient context for its reasoning process, our Translator module initially generates a complete and comprehensive set of textual conditions.
+However, to mitigate information inconsistency across modalities in the final synthesized dataset, the Translator performs a crucial refinement step.
+It systematically prunes any explicit constraint-based information that is visually rendered in the image (e.g., numerical values for segment lengths or angles).
+This process prevents the model from developing “shortcuts” based on redundant textual data, which would impair its cross-modal reasoning capabilities.
+The final text exclusively retains abstract information about entities (e.g., ”isosceles triangle”) and their semantic relations (e.g., ”point O is the incenter of triangle ABC”), compelling the model to ground its understanding in the visual domain.
+A.4 Reasoner
+We employ DeepSeek-R1 as our reasoning LLM to conduct a forward search process.
+At each step, the model receives the conditions of the current state and is prompted to infer a single, new contingent and computable geometric property or relationship that augments this state.
+The specific prompt used for this task is shown in Box A.4.
+The prompt for Reasoner to reverse search Use the mathematics you know to make simple and accurate inferences and get the conclusions based on image descriptions.
+Focus on extracting as much information as possible.
+Keep the final answer of all reasoning to one decimal place.
+Use <<conclusion>> <</conclusion>> to include all your inferences.
+Example is as follows: Output: <<conclusion>> 1.
+The area of Triangle SFD is 18.0.
+2.
+The length of line EF is 3.9.
+3.
+The measure of angle DFS is 29.4°.
+<</conclusion>> Here is the image description input: Current state
+A.5 Verifier
+We employ DeepSeek-V3 to conduct the forward verification process. The prompt used for this task is detailed in Box A.5.
+The prompt for the Verifier to Forward Validation Here is a geometry problem with the following information: Image description: Initial state Question: Questions from Reasoner Following the above condition and question, think step by step and answer the following question directly.
+After thinking process, you should provide your final concise reasoning steps in <<steps>><</steps>> tags and your final answer in <<answer>> </</answer>> tags.
+The answer to the question is only allowed to contain one number or angle.
+Appendix B More Details of Datasets
+B.1 Statistics of Datasets
+Detailed numerical statistics and element distribution for the NeSyGeo-Caption and NeSyGeo-CoT datasets are presented in Table 7 and  7.
 For element distribution statistics, we randomly sampled 1.8k Geo-DSL sequences corresponding to images from each dataset, counting the frequency of different geometric elements.
@@
 Figure 6: Frequency of different geometric elements. To facilitate interpretation, these elements are converted into corresponding natural language descriptions.
-Appendix C Additional Experimental Details and Results
-We utilized the VLM-R1 [22] framework for RL experiments, conducted on 6 vGPU-32 GB.
+B.2 Difficulty Definitions
+To support evaluation and training paradigms such as curriculum learning, we annotate each sample with a difficulty level.
+Given that geometric reasoning tasks primarily require models’ image perception and logical reasoning capabilities, we scientifically define the difficulty level as
+$\alpha\times\text{perception difficulty}+(1-\alpha)\times\text{reasoning difficulty},$ (3)
+where perception difficulty is the number of Geo-DSL statements, and reasoning difficulty is the number of reasoning steps. Here, we set alpha as 0.3.
+Each synthesized sample includes detailed meta-information stored as a symbolic form based on our Geo-DSL language.
+This symbolic form accurately describes the geometric setup and offers promising directions for future research.
+For instance, valid geometric configurations could be generated by augmenting or mutating existing symbolic forms within constrained parametric bounds.
+B.3 Details of NeSyGeo-Test Benchmark.
+Our NeSyGeo-Test benchmark comprises 2668 Q&A pairs.
+Consistent with the training set, numerical annotations are embedded in the image space, with only essential conditions and questions provided in the text.
+The type of numerical quantity categorizes the dataset sought: Angle (658 pairs), Shape (730 pairs), and Length (1280 pairs).
+Shape type includes shape area and perimeter, while length type includes edge and arc lengths.
+Based on the difficulty level in  B.2, problem difficulty is divided into three levels: Easy (1537 pairs), Medium (908 pairs), and Hard (223 pairs).
+Appendix C More Details and Results of Experiments
+C.1 Settings of Dataset Generation
+To synthesize a diverse dataset, we set hyperparameters for generating images by configuring the step count, length range, and angle range.
+The step count $N$ ranges from one to four.
+The line length is defined within the basic range $l_{\text{min}},l_{\text{max}}]=[1,5]$, scalable by any multiple of 2 (e.g., $[4,20]$).
+The angle is constrained to multiples of 15° within $15^{\circ},165^{\circ}]$, with increased weights assigned to special angles.
+The temperature hyperparameters $\tau_{e}$ and $\tau_{a}$ were set to a default value of 1.
+We generate various types of weight matrices $A$ and $I$ by adjusting their corresponding values.
+C.2 Human Preference evaluation
+To rigorously assess the quality of our generated data, we conducted a comprehensive human evaluation study.
+We compared our NeSyGeo-CoT dataset against two prominent, automatically generated geometry datasets: G-LLaVA and MAVIS.
+Methodology
+We recruited 40 volunteers.
+To ensure a proficient and consistent level of domain expertise, a prerequisite for participation was holding a university undergraduate degree or higher.
+To ensure the objectivity and accuracy of the evaluation, participants were informed that the study pertained to middle school mathematics education and that their primary task was to meticulously verify the correctness of the provided solutions and final answers for each problem.
+We guaranteed participants that their responses would be completely anonymized to protect their privacy, with the right to withdraw at any time without penalty.
+Each participant was presented with a questionnaire containing nine geometric problems.
+To mitigate potential ordering effects and sampling bias, the problems were composed of three samples randomly and independently drawn from each of the three datasets being evaluated.
+Evaluation Criteria
+Participants were asked to rate each sample on a 5-point Likert scale across five key dimensions designed to capture the multifaceted quality of a multimodal reasoning problem:
+Image
+This dimension evaluates the visual output.
+Image Correctness assesses whether the image 1)accurately reflects all entities and constraints from the text premise.2) has no conflict itself Image Appearance rates the human aesthetic quality, including the clarity of annotations and overall layout.
+Question
+This assesses the generated question.
+Question Correctness evaluates its logical validity and solvability based on the provided context, while Question Fluency rates its grammatical structure and naturalness.
+Reasoning
+This pertains to the generated solution.
+Reasoning Correctness scrutinizes the logical and mathematical validity of the step-by-step thought process.
+Reasoning Fluency assesses the clarity and readability of the explanation.
+Overall
+This captures the evaluators’ subjective impressions.
+Overall Reality gauges the plausibility of the problem, assessing if it feels like a genuine question from a textbook or exam.
+Overall Difficulty is the participant’s rating of the cognitive challenge involved.
+Answer
+Answer Correctness provides the definitive assessment of problem-solving capability by verifying if the final numerical answer derived from the reasoning process is objectively correct.
+Results The aggregated results are presented in Picture 4.
+The findings indicate a clear preference for our NeSyGeo-CoT dataset.
+While all datasets scored perfectly on answer correctness—an expected outcome for synthetically generated data—our dataset was rated significantly higher in Reasoning Path Quality and Overall Difficulty.
+This suggests that our framework not only generates valid and high-fidelity problems but also produces content that is more complex and challenging, thereby validating its effectiveness for advancing the reasoning capabilities of MLLMs.
+C.3 Reinforcement Learning
+We utilized the VLM-R1 (Shen et al., 2025) framework for RL experiments, conducted on 6 vGPU-32 GB.
 We set epochs to 2, num generations to 6, batchsize to 1.
 To enhance the visual perception capabilities of MLLMs, parameters of the language model and vision modules are set to be trainable.
+Table 8 presents the detailed performance of models trained on various automatically synthesized datasets across the MathVerse benchmark.
+Models trained using our dataset demonstrate superior performance on most metrics compared to others, exhibiting substantial performance gains relative to the base model.As shown in Figure 7.
+We also evaluated model performance as RL training steps increased when using NeSyGeo-CoT.
+Most metrics improved with more training steps, demonstrating the robustness and effectiveness of our datasets.
 Table 8: Detailed RL experiments evaluation on MathVerse. Here, ‘AGL’, ‘ARA’, ‘LTH’, and ‘PG’ denote angle, area, length, and plane geometry, respectively.
@@
 InternVL2.5-4B+RL 45.6 20.9 37.9 40.2 45.1 26.4 36.2 38.2 29.0 27.5 34.1 31.8
-For SFT experiments, we employed the LLaMA-Factory [40] framework on 2 A800 GPUs with LoRA.
+Figure 7: Model performance on Mathverse as the RL training steps increase.
+With InternVL2.5-4B as our base model, most metrics exhibit progressive improvement throughout training, demonstrating the robustness and effectiveness of our datasets.
+C.4 Supervised Finetuning
+For SFT experiments, we employed the LLaMA-Factory (Zheng et al., 2024) framework on 2 A800 GPUs with LoRA.
 We set the learning rate of $1\times 10^{-5}$, LoRA rank of 64, and use Adam optimization.
 Training on NeSyGeo-Caption used 1 epoch, while NeSyGeo-CoT used 2 epochs.
-Table 7 presents the detailed performance of models trained on various automatically synthesized datasets across the MathVerse benchmark.
-Models trained using our dataset demonstrate superior performance on most metrics compared to others, exhibiting substantial performance gains relative to the base model.As shown in Figure 9.
-We also evaluated model performance as RL training steps increased when using NeSyGeo-CoT.
-Most metrics improved with more training steps, demonstrating the robustness and effectiveness of our datasets.
-To investigate visual diversity directly across datasets, we randomly sampled 1k images from each, extracted features using ResNet, and visualized them via t-SNE.
-As illustrated in Figure 10, G-LLaVA—having augmented only the textual components of its base dataset—displays a distinctly non-uniform distribution in the image feature space.
-Conversely, our method exhibits uniform feature distributions, underscoring the substantial visual diversity of images generated by our approach.
-Figure 9: Model performance on Mathverse as the RL training steps increase.
-With InternVL2.5-4B as our base model, most metrics exhibit progressive improvement throughout training, demonstrating the robustness and effectiveness of our datasets.
-Figure 10: T-SNE of the image features of different automatic frameworks.
-Our datasets exhibit uniform feature distributions, underscoring the substantial visual diversity of images generated by the NeSyGeo framework.
-Appendix D More Examples of NeSyGeo-CoT Dataset
-Figure 11: Examples of the NeSyGeo-CoT dataset.
-Each sample comprises a symbolic image definition based on our Geo-DSL language, a high-quality annotated image, a concise text caption, diverse Q&A pairs, and a detailed reasoning process step-by-step.
-We present more examples from the NeSyGeo-CoT dataset in Figure 11.
-Our bidirectional conversion engine can generate high-quality visual images from a symbolic form based on our Geo-DSL language.
-To further enhance image diversity, the engine introduces variability by randomly selecting values for the unit length and applying random rotations to the generated diagrams during creation.
-While other visual attributes, such as element and background colours, could also be randomized, they were set to default values in our current synthesis process.
-Our symbolic language helps identify parts of the image, and our conversion process ensures the images are geometrically correct.
-Due to LLMs’ powerful search and reasoning capabilities, we obtain diverse Q&A pairs concerning properties like lengths, angles, or areas alongside high-quality CoT step-by-step.
-This process, involving backwards search across the geometric space defined by the symbolic form and forward validation, ensures the correctness of the numerical answers, thereby enriching textual diversity.
-To support evaluation and training paradigms such as curriculum learning, we annotate each sample with a difficulty level.
-Given that geometric reasoning tasks primarily require models’ image perception and logical reasoning capabilities, we scientifically define the difficulty level as
-[…]
-where perception difficulty is the number of Geo-DSL statements, and reasoning difficulty is the number of reasoning steps.
-Each synthesized sample includes detailed meta-information stored as a symbolic form based on our Geo-DSL language.
-This symbolic form accurately describes the geometric setup and offers promising directions for future research.
-For instance, valid geometric configurations could be generated by augmenting or mutating existing symbolic forms within constrained parametric bounds.
-Appendix E Details of NeSyGeo-Test Benchmark.
-Our NeSyGeo-Test benchmark comprises 2668 Q&A pairs.
-Consistent with the training set, numerical annotations are embedded in the image space, with only essential conditions and questions provided in the text.
-The type of numerical quantity categorizes the dataset sought: Angle (658 pairs), Shape (730 pairs), and Length (1280 pairs).
-Shape type includes shape area and perimeter, while length type includes edge and arc lengths.
-Based on the difficulty level in  D, problem difficulty is divided into three levels: Easy (1537 pairs), Medium (908 pairs), and Hard (223 pairs).
+C.5 Diversity of Dataset
+Diversity: Datasets synthesized by our NeSyGeo framework exhibit high diversity in text and visual features.
+A critical challenge for automatic data synthesis methods is whether the dataset is sufficiently diverse to avoid quality degradation due to potential overfitting risks from inherent domain constraints.
+We employ t-SNE (van der Maaten & Hinton, 2008) dimensionality reduction for mapping in text space to evaluate the diversity across different methods.
+This analysis allows us to assess the diversity of the textual descriptions themselves.
+Given that in geometric problems, the text conditions depict specific visual elements, and the diversity observed in the text space also serves as a valuable indicator of the diversity in the corresponding visual diagrams.
+To ensure a fair comparison, we remove all prompts related to guiding large models, retaining only condition and question texts, and randomly sample 5k texts from each dataset.
+The results are illustrated in Figure 8.
+Our method and G-LLaVA (Gao et al., 2025) exhibit uniformly distributed features in the space, indicating low data overlap and high diversity.
+In contrast, R-CoT and MAVIS display varying degrees of clustered distribution, indicating more feature-similar samples.
+Figure 8: T-SNE of the text features of different automatic frameworks.
+The G-LLaVA method augments the text space on the manually annotated GeoQA dataset.
+Thus, its text diversity can approximate that of real data more closely.
+Similar to G-LLaVA, our method exhibits a uniform distribution in the space, demonstrating superior diversity.
+Appendix D NeSyGeo-Test Benchmark
+To probe the genuine cross-modal reasoning capabilities of current models, our empirical analysis is grounded in the NeSyGeo-test benchmark.
+This dataset was engineered to address the shortcomings of its predecessors.
+It comprises high-fidelity, programmatically rendered images and mitigates modality imbalance by ensuring crucial information is distributed across both the visual and textual inputs.
+A cornerstone of its design is the elimination of textual redundancies, a feature that compels models to ground their deductions in the visual context and prevents reliance on unimodal heuristics.
 Evaluation results on current mainstream open-source and closed-source MLLMs are shown in Table 9.
+D.1 Quantitative Performance
+A key observation is the benchmark’s considerable difficulty, which challenges even state-of-the-art MLLMs.
+This is particularly evident in the Hard difficulty tier, where the highest accuracy achieved by any closed-source model is merely 56.7%, while the leading open-source model, Qwen2.5-VL-32B-Instruct, scores only 43.3%.
+These results underscore the significant room for improvement in complex geometric reasoning.
+Furthermore, the effectiveness of our difficulty stratification is validated by the consistent performance degradation exhibited by most models as they progress from Easy to Medium and Hard questions.
+This trend confirms the rational design of our benchmark’s difficulty levels.
+Finally, the results reveal a clear performance gap between model types, with leading closed-source models like Claude-3.5-Sonnet (74.5%) consistently outperforming the best open-source models (55.8%), highlighting the current advantages of proprietary systems on such specialized tasks.
 Table 9: NeSyGeo-Test Benchmark on several mainstream MLLMs. The highest accuracy for open-source and closed-source MLLMs is marked in red and blue, respectively.
@@
 Qwen2.5-VL-3B-Instruct 3B 44.1 31.8 27.8 30.5 29.1 31.9 30.9
-Qwen2.5-VL-7B-Instruct 7B 38.3 30.7 32.2 36.8 27.5 32.3 43.3
-InternVL2.5-4B 4B 53.8 53.1 53.5 62.9 49.4 32.0 53.4
-InternVL2.5-8B 8B 55.9 56.2 55.4 64.5 49.4 43.3 55.8
-LLaVA-NeXT-7B 7B 23.7 15.5 15.4 18.6 14.7 13.3 6.4
-LLaVA-NeXT-13B 13B 15.7 15.6 16.0 15.8 16.0 15.2 15.8
+Qwen2.5-VL-7B-Instruct 7B 38.3 30.7 32.2 36.8 27.5 32.3 33.3
+Qwen2.5-VL-32B-Instruct 7B 38.3 30.7 32.2 65.8 49.4 41.7 58.2
+InternVL2.5-8B 8B 55.9 56.2 55.4 64.5 49.3 43.3 55.8
+LLaVA-NeXT-7B 13B 15.7 15.6 16.0 15.8 16.0 15.2 15.8
+LLaVA-NeXT-13B 7B 23.7 15.5 15.4 18.6 14.7 13.3 16.4
 LLaVA-NeXT-34B 34B 23.7 21.0 18.5 19.1 21.5 19.2 19.9
@@
 Gemini-2.0-Flash 36.8 60.4 67.7 54.3 63.8 61.0 58.1
-Appendix F Limitations
-  • Limited Training Paradigm: Our current evaluation of the NeSyGeo dataset relies on a simple training paradigm to assess its efficacy for automated data generation.
-This approach lacks advanced training strategies, such as CLIP alignment or curriculum learning, which restricts the development of specialized models optimized for geometric reasoning tasks.
-  • Restricted Domain Scope: The NeSyGeo framework is currently tailored to plane geometry, limiting its generalizability to other domains.
-However, we believe that for multimodal datasets in other domains, we can similarly achieve synthesis by defining symbolic statements, shifting the synthesis process to a controllable symbolic space, and constructing a symbolic-to-image engine, which we plan to explore in our future work.
-  • Dependency on External APIs: The construction of Q&A pairs in this study partly relies on the reasoning capabilities of LLMs.
-This dependency increases generation costs and introduces potential inconsistencies.
-We aim to develop an automated solver that conducts search and validation directly within the symbolic space, thereby removing reliance on LLMs.
-This could further reduce costs and ensure complete rigor.
-Appendix G Details of Prompts in Reverse Search and Forward Validation
-In our automatic synthesis framework, we employ DeepSeek R1 as the expert LLM for reverse search and DeepSeek V3 for forward validation.
-The specific prompts utilized are detailed in Figures 12 and 13, respectively.
-Note that the blue text in these prompts is substituted with actual content.
-Appendix H Detailed Definition of Geo-DSL
-Geo-DSL adopts an entity-relation-constraint framework to define geometric elements in plane geometry, encompassing 13 types of points, 7 types of lines, 3 types of angles, and 14 types of shapes.
-Representative examples of symbolic statements and their corresponding natural language descriptions are illustrated in Figures 18, 15, 16, and 17.
+D.2 Analysis of Failure Cases
+Despite recent advancements, large multimodal models (LMMs) still exhibit several fundamental failure modes when tasked with complex geometric reasoning.
+Through our analysis, we identify and categorize three primary types of error:
+Semantic Grounding Errors: This class of error occurs when the model fails to correctly associate visual annotations with their corresponding geometric entities.
+For instance, a model might misattribute the numerical value of a specific angle to an adjacent, unlabeled angle.
+While the model may occasionally identify such a contradiction through subsequent deductive steps, this process is unreliable.
+A more robust solution involves enhancing the model’s architecture to support iterative self-correction, allowing it to perform consistency checks across its reasoning chain and rectify initial grounding errors.
+Deficiencies in Latent Property Inference: Models often overlook implicit geometric properties that are not explicitly annotated in the diagram, thereby increasing the perceived complexity of the problem.
+A common example is the failure to infer the congruence of sides in an isosceles triangle, which leads the model into a spurious exploration of the solution space through exhaustive, case-based analysis.
+This frequently results in wasted computational budget on incorrect deductive paths, sometimes causing premature termination due to exceeding context length limits.
+Future work should focus on improving the model’s perceptual acuity and multi-step reasoning capabilities, enabling it to prune invalid branches of logic efficiently, even when faced with incomplete information.
+Vulnerability to Deductive Fallacies: The reasoning process can be derailed by irrelevant visual information or distractors within the image, causing the model to commit logical errors.
+This often results in the generation of self-contradictory statements or conclusions that are tangential to the problem’s solution path.
+To mitigate this, we propose the integration of a formal verification module.
+Such a neuro-symbolic validator would serve as a safeguard, programmatically ensuring that each step in the chain-of-thought is a logically sound consequence of the established premises and previously derived facts, thereby enforcing a more rigorous deductive process.
+We present additional failure cases of Gemini-2.0-Flash in Figures 9 and 10, where the error types made by the model during the geometric reasoning task are highlighted with corresponding colors.
+Figure 9: Failure cases of NeSyGeo-Test.
+Figure 10: Failure cases of NeSyGeo-Test.
+D.3 Future Direction
+Addressing these challenges points toward several key research directions for advancing multimodal geometric problem-solving.
+The path forward likely involves a tighter integration of perceptual modules with symbolic reasoning engines, moving toward more capable architectures.
+Furthermore, progress will depend on the development of architectures that explicitly support stateful, long-horizon reasoning, allowing them to maintain and update a coherent ”mental model” of the geometric state.
+Finally, enabling models to achieve self-improvement through reinforcement learning loops, where rewards are provided by feedback from formal verifiers or successful problem completion, represents a crucial frontier for creating more robust and reliable geometric reasoning agents.
+Appendix E Case Analysis
+E.1 Comparison with other Popular Geometry Datasets
+To facilitate comparison of dataset characteristics synthesized by our method and other popular approaches, we showcase a randomly selected example from NeSyGeo-CoT alongside each of the different approaches in Figure 11.
+Geometry-3K is a manually synthesized dataset, while the remaining approaches employ automatic generation techniques.
+To ensure a fair comparison, we standardize the text format by removing model-guiding prompts and appending options when present.
+Furthermore, we annotate each sample with image pixels and CoT word counts.
+Compared to other datasets, our dataset features clear, human-aesthetically pleasing images, high-quality step-by-step reasoning chains, symbolic form meta-information enabling subsequent image augmentation and mutation, and well-distributed conditional information between images and text.
+Additional examples of our NeSyGeo-CoT dataset can be found in the Appendix 12.
+Figure 11: Comparison of NeSyGeo-CoT dataset with other Popular Geometry Datasets.
+Geometry-3K is a manually synthesized dataset, while the remaining approaches employ automatic generation techniques.
+Our dataset features clear, human-aesthetically pleasing images, high-quality step-by-step reasoning chains, symbolic form meta-information enabling subsequent image augmentation and mutation, and well-distributed conditional information between images and text.
+E.2 More examples of our datasets
+We present more examples from the NeSyGeo-CoT dataset in Figure 12.
+Our bidirectional conversion engine can generate high-quality visual images from a symbolic form based on our Geo-DSL language.
+To further enhance image diversity, the engine introduces variability by randomly selecting values for the unit length and applying random rotations to the generated diagrams during creation.
+While other visual attributes, such as element and background colours, could also be randomized, they were set to default values in our current synthesis process.
+Our symbolic language helps identify parts of the image, and our conversion process ensures the images are geometrically correct.
+Figure 12: Examples of the NeSyGeo-CoT dataset.
+Each sample comprises a symbolic image definition based on our Geo-DSL language, a high-quality annotated image, a concise text caption, diverse Q&A pairs, and a detailed reasoning process step-by-step.
+E.3 Case of Reasoner and Verifier
+Figure 13: Case of reasoner and verifier.
+Figure 14: Cases filtered out due to an inconsistency detected during the verification phase.
+To provide a transparent illustration of our methodology, we present a concrete case study in Figure 13.
+This example details the step-by-step reverse search conducted by the Reasoner and the subsequent forward reasoning performed by the Verifier.
+Furthermore, to highlight the effectiveness of our validation mechanism, we also showcase an instance in Figure13that was filtered out due to an inconsistency detected during the verification phase.
+Appendix F Use of LLMs
+Within the Reasoner and Verifier modules of NeSyGeo, we employ LLMs to perform deductive geometric reasoning.
+This is actualized through a two-stage generation process, comprising a step-by-step reverse search and a reliable forward validation.
+The formidable reasoning capabilities and extensive pre-trained knowledge inherent in LLMs facilitate the efficient expansion of reasoning states, while simultaneously ensuring the fidelity and diversity of the resulting reasoning paths.
+All technical contributions, including the design of experiments, comparison of MLLMs (e.g., Qwen and GPT-4o), and interpretation of results, were fully conceived and conducted by the authors.
+The authors take full responsibility for the correctness, originality, and integrity of all scientific content presented in the paper.
+Appendix G Future Work
+We intend to extend NeSyGeo to other multimodal domains, such as analytical geometry and visual question answering.
+This extensibility will be achieved by defining new domain-specific languages, corresponding synthesis rules within the symbolic space, and tailored conversion engines.
+Furthermore, we plan to develop an automated symbolic solver capable of conducting search and validation directly within the symbolic space.
+This would remove reliance on LLMs, potentially reducing generation costs and ensuring complete correctness of the datasets.
+Appendix H Details of Domain-Specific Languages
+H.1 Definition of Geo-DSL
+Geo-DSL adopts an entity–attributes-relations framework to define geometric elements in plane geometry, encompassing 13 types of points, 7 types of lines, 3 types of angles, and 14 types of shapes.
+Representative examples of symbolic statements and their corresponding natural language descriptions are illustrated in Figures 19, 16, 17, and 18.
 With a single statement, Geo-DSL uniquely specifies spatial elements, ensuring the accuracy of geometric synthesis while significantly facilitating parsing and transformation by our conversion engine.
@@
 These advantages make Geo-DSL a robust and versatile solution for generating high-quality multimodal geometric reasoning data.
-Appendix I Detailed Actions in Symbolic Spaces
+H.2 Definition of Actions in Symbolic Sequence Generation
 As illustrated in Figure 15, we enumerate all statements within the action space defined by our Geo-DSL. The content within square brackets denotes annotations for each statement.
@@
 For different types of actions, we provide a concrete example for each, highlighted with a gray background to indicate the available action space when the respective element is selected.
-Figure 12: Prompt for Deepseek R1 in reverse search.
-Figure 13: Prompt for Deepseek V3 in forward validation.
 Figure 15: Detailed Actions in Symbolic Spaces. Actions can be categorized into four parts based on the type of selected geometric element: line-based, point-based, shape-based, and angle-based.